### v1

# Enhanced Generative Recommendation via Content and Collaboration Integration

Abstract.
Generative recommendation has emerged as a promising paradigm aimed at augmenting recommender systems with recent advancements in generative artificial intelligence. This task has been formulated as a sequence-to-sequence generation process, wherein the input sequence encompasses data pertaining to the user’s previously interacted items, and the output sequence denotes the generative identifier for the suggested item. However, existing generative recommendation approaches still encounter challenges in (i) effectively integrating user-item collaborative signals and item content information within a unified generative framework, and (ii) executing an efficient alignment between content information and collaborative signals.
In this paper, we introduce content-based colla-borative generation for recommender systems, denoted as ColaRec. To capture collaborative signals, the generative item identifiers are derived from a pretrained collaborative filtering model, while the user is represented through the aggregation of interacted items’ content. Subsequently, the aggregated textual description of items is fed into a language model to encapsulate content information. This integration enables ColaRec to amalgamate collaborative signals and content information within an end-to-end framework. Regarding the alignment, we propose an item indexing task to facilitate the mapping between the content-based semantic space and the interaction-based collaborative space. Additionally, a contrastive loss is introduced to ensure that items with similar collaborative GIDs possess comparable content representations, thereby enhancing alignment. To validate the efficacy of ColaRec, we conduct experiments on three benchmark datasets. Empirical results substantiate the superior performance of ColaRec.
Recommender System, Generative Recommendation, Collaborative Filtering, Generative Model, Language Model
isbn: ccs: Information systems Recommender systemsccs: Information systems Retrieval models and rankingccs: Information systems Novelty in information retrieval
1. introduction
Recommender systems are widely deployed to discover user interests and provide personalized information services (Hu et al., 2018; Qi et al., 2020; Yuan et al., 2019). Recently, generative models, such as large language models (LLMs) (Zhao et al., 2023; Chang et al., 2023) and diffusion models (Yang et al., 2023; Cao et al., 2022), have gained prominence in advancing artificial intelligence. In such a context, generative recommendation (Wang et al., 2023b; Bao et al., 2023; Si et al., 2023; Rajput et al., 2023; Geng et al., 2022; Ji et al., 2023) has emerged as a new paradigm to enhance recommendation agents with end-to-end generation capability of these models.
Generative Recommendation. Generative recommendation assigns each item with a unique sequence of tokens as the item’s generative identifier, a.k.a., GID, then the recommendation is formulated as a sequence-to-sequence generation task, where the input is historical user-item interactions while the output sequence refers to the GID of the recommended item (Rajput et al., 2023; Si et al., 2023; Hua et al., 2023) [^1] Wang et al. (2023a) proposed another paradigm to directly generate new content, e.g., images, for recommendation. However, their methods are tailored for the generation of virtual content and cannot be used for recommendation of concrete items, such as the item recommendation in an e-commerce platform. In this paper, we target on the recommendation of concrete items..
Figure 1. Comparison between conventional itemIDs and GIDs. GIDs contain more concrete correlations.
Compared with conventional itemIDs with an assigned single random token, the sequential tokens of a GID contain more explicit representation information. As shown in Figure 1, we see correlated GIDs denote correlated items, and thus help the recommender to gain better performance. Generative recommendation provides an end-to-end paradigm to effectively utilize generative models for recommendation. Tay et al. (2022) initially introduced Differentiable Search Indexing (DSI), which marks an early endeavor in employing generative models for information retrieval. DSI has inspired a series of studies on generative recommendation (Rajput et al., 2023; Si et al., 2023; Petrov and Macdonald, 2023). The promising results of these works demonstrate the potential of generative recommendation.
Limitations of Current Approaches. Existing generative recommendation approaches still have the following two limitations, leading to suboptimal performance.
Firstly, existing methods fail to effectively model collaborative signals and item content information in a unified framework. Collaborative signals refer to the knowledge contained in the user-item interactions while item content information refers to the textual description of items, as shown in Figure 2. On one hand, several methods utilize either the item title or hierarchical content embeddings of items to construct GIDs (Li et al., 2023c; Rajput et al., 2023). However, these methods only consider the item content information while the collaborative signals between users and items are overlooked. On the other hand, Si et al. (2023) construct GIDs using item embeddings of a pretrained SASRec (Kang and McAuley, 2018). Although the GID in this work contains the item-item sequential (collaborative) connections, the proposed recommendation framework in their research fails to model the item content information. To achieve satisfying recommendation results, we argue that the generative recommenders should be able to jointly model the item content information and the user-item collaborative signals in a unified generative framework.
Secondly, existing generative recommendation methods cannot effectively align item content information and collaborative signals. Although Hua et al. (2023) proposed to construct the GID using the concatenation of a semantic string learnt from the item content and a sequence of tokens learnt from the item-item co-occurrence matrix, the naive concatenation without a learning process is not effective to perform the alignment, leading to sub-optimal performance. The rich content information can assist the recommendation models in obtaining more informative and comprehensive item representations, thus enhancing the modeling of user-item collaborative interactions, and vice visa. To this end, the alignment between collaborative signals and content information is fundamental for further enhancing generative recommenders. However, we argue that a successful alignment should be achieved through an explicit learning process, e.g.,the mapping between the content-based semantic space and the interaction-based collaborative space.
Figure 2. Illustration of collaborative signals and content information. Collaborative signals refer to the knowledge contained in user-item interactions while content information refers to the textual description of items.
Proposed Methods. In this paper, we propose content-based colla-borative generation for recommender systems (ColaRec), a generative recommendation model which unifies both item content information and user-item collaborative interaction signals in a sequence-to-sequence generation framework. Taking a particular user as an example, the input sequence of ColaRec consists of unordered [^2] In this paper, we target on the general recommendation task other than sequential recommendation. To this end, we use unordered item tuples as the input. tuples with each tuple describing the content information of one interacted item of this user. Then an encoder-decoder framework is used to generate the GID of the target item.
To address the first limitation, we propose two tailored designs to jointly model user-item collaborative signals and item content information. Firstly, we propose to construct GIDs using item representations obtained from a pretrained collaborative filtering model. In this paper, we use LightGCN (He et al., 2020) as the pretrained model. The LightGCN model is trained on the user-item interaction graph and thus the constructed GID can effectively encodes the user-item collaborative signals. Note that the LightGCN model can also be alternated with other models. Secondly, in ColaRec, the user is represented as the content aggregation of tuples with each tuple describing one her/his historically interacted item. This design also keeps inline with the nature of collaborative filtering. The aggregation of content-based tuples is fed to an encoder-decoder based language model to effectively capture the textual content information. In this paper, we use a fine-tuned T5 (Raffel et al., 2020) as the involved language model. We leave the investigation of larger language models as one of our future works.
To address the second limitation, we propose an auxiliary item indexing task which targets on mapping the item side information into the GID of this item through the same encoder-decoder model. More precisely, the item side information contains both the textual content information and a set of users who have interacted with this item. To this end, the indexing task maps both the item content information and user-item interaction signals into the constructed GID, achieving better alignment between item content information and user-item collaborative signals. Besides, we further propose a contrastive loss to ensure that items with similar collaborative GIDs are also similar in the content-based semantic space. Such the contrastive loss also helps to perform alignment between content information and collaborative signals.
To demonstrate the effectiveness of the proposed ColaRec, we conduct extensive experiments on three public accessible benchmark datasets. Experimental results show that the proposed ColaRec outperforms related state-of-the-art baselines.
Main Contributions. Our main contributions are as follows:
  • We propose ColaRec, an end-to-end generative recommender which utilizes an encoder-decoder based language model to jointly model item content information and user-item collaborative signals for recommendation.
  • We propose an auxiliary item indexing task and a contrastive loss to perform better alignment between item content information and user-item collaborative signals to further enhance the performance of generative recommendation.
  • We conduct extensive experiments on three datasets to demonstrate the effectiveness of the proposed ColaRec. Experimental results show superior recommendation performance of ColaRec.
2. related work
In this section, we review related prior literature on collaborative filtering and generative models.
2.1. Collaborative Filtering
Collaborative filtering (CF) is one of the most representative methods to build a recommendation agent. CF believes that a user can be represented as the aggregation of her/his interacted items, and vice visa. The keystone to conduct collaborative filtering is the user-item interaction matrix. Early approaches (Koren et al., 2009; Rendle et al., 2012) are based on matrix factorization (MF) to jointly model the latent space for users and items. Due to the expressiveness of deep neural networks, plenty of research (He et al., 2017; Mao et al., 2021; Zheng et al., 2017; Chen et al., 2018; Gao et al., 2020) has been conducted to enhance CF through deep learning. Besides, since the user-item interaction signals can be naturally encoded into an interaction graph, graph neural networks (GNN) also shed lights in the field of CF. Berg et al. (2017) proposed to use graph convolution for matrix completion. Wang et al. (2019) proposed the NGCF model to use GNN for CF. He et al. (2020) proposed the LightGCN model to simplify GNN for recommender systems, leading to a simpler and linear CF model. LightGCN serves as one of the most popular GNN-based CF approaches due to its effectiveness. Lin et al. (2022) proposed NCL to introduce contrastive learning into graph-based CF. Besides, content information has also been utilized to enhance the CF model. Rendle (2010) proposed the notable factorization machine (FM) to extend MF for categorical contextual features. He and McAuley (2016); Wei et al. (2019, 2020); Wu et al. (2017) proposed to enhance CF with visual or text features. Li et al. (2023b) proposed RecFormer to model long text sequences for recommendation.
Different with existing CF methods, in this paper we focus on the new paradigm of generative recommendation, where the recommended item is generated through the decoder of a sequence-to-sequence model. Generative recommendation provides new potential to utilize advances in the field of generative artificial intelligence to enhance the recommendation task.
2.2. Generative Models
Generative models have become a hot research topic to generate new content, such as images and text, from the collected data. Variational Autoencoders (VAEs) (Kingma and Welling, 2014; Higgins et al., 2016) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Creswell et al., 2018), are two kinds of representative generative models. VAEs aim to learn a probabilistic mapping from the observed data to a latent space, and then leverage variational inference to approximate the posterior distribution of latent variables. GANs introduce a min-max gaming approach with a generator attempting to generate simulated samples, and the discriminator distinguishing between real and generated data. Recently, Transformers (Vaswani et al., 2017) have shown promise in language generation, leading to notable LLMs like GPTs (Generative Pre-trained Transformers). Besides, diffusion models (Ho et al., 2020; Yang et al., 2023; Cao et al., 2022) have also shown promising results in content generation.
Generative models for document retrieval. Generative retrieval has drawn increasing attention to conduct document retrieval with a sequence-to-sequence framework, where the input sequence is the query and the output sequence is the document identifier. The bloom of LLMs encourages generative retrieval to better understand semantic patterns stored in the retrieval process. Tay et al. (2022) firstly proposed a differentiable search index (DSI) for document retrieval, which assigns each document with a structured semantic meaningful identifier. DSI can be regarded as the early attempt to conduct generative retrieval. Plenty of work has been conducted following this research line. For example, strings including titles (Tang et al., 2023; Cao et al., 2021), n-grams (Bevilacqua et al., 2022; Chen et al., 2023) or URLs (Zhou et al., 2022) are explored to construct the document identifier. Wang et al. (2022a) proposed NCI which produces a hierarchical document identifier to get fine-grained semantic mapping, and a novel decoder to consider the prefixes of document identifiers. GenRet (Sun et al., 2023) learns to tokenize the document into a short discrete representation via a discrete auto-encoding approach, and utilizes progressive training and diverse clustering to assist the model training process.
Figure 3. Overview of ColaRec. ColaRec assigns each item with a GID obtained from a GNN-based CF model. ColaRec consists two tasks. User-item recommendation aims to map the user’s interacted items with textual content into the GID of the recommended item, i.e., $\mathcal{L}_{\text{rec}}$. Item-item indexing targets on the mapping from item side information into the item’s GID, i.e., $\mathcal{L}_{\text{index}}$. Besides, a ranking loss $\mathcal{L}_{\text{bpr}}$ and a contrastive loss $\mathcal{L}_{\text{c}}$ are also introduced. A tuned T5 is used to accomplish these tasks.
Generative models for recommendation. Generative models have also been utilized for recommender systems. Early works focus on utilizing VAEs (Liang et al., 2018; Shenbin et al., 2020; Cai and Cai, 2022) or GANs (He et al., 2018; Guo et al., 2022; Wang et al., 2022b) for recommendation. Recently, diffusion models have also been utilized to perform item recommendation (Wang et al., 2023b; Li et al., 2023a). Besides, LLM-based recommendation has also been proposed (Bao et al., 2023; Zhang et al., 2023; Petrov and Macdonald, 2023). These works explore the ability of LLMs for recommendation through parameter efficient tuning (Bao et al., 2023) or instruction tuning (Zhang et al., 2023). Meanwhile, several works attempt to convert various recommendation tasks into a unified natural language generation task, and train recommenders with multi-task optimization, like P5 (Geng et al., 2022) and M6-Rec (Cui et al., 2022).
Notably, TIGER (Rajput et al., 2023) is a representative generative recommendation method which uses an RQ-VAE (Zeghidour et al., 2022) to construct the GID and then uses encoder-decoder based transformers to generate sequential recommendation. Si et al. (2023) proposed to construct the GID from a pretrained SASRec (Kang and McAuley, 2018) model for sequential recommendation. Hua et al. (2023) further investigated the effect of item identifier construction for generative recommendation.
Despite the arise of generative recommenders, existing methods still suffer from the ineffective infusion of collaborative signals and item content information. How to jointly model and align collaborative signals and content information in a sequence-to-sequence generative framework for recommender systems is still an open research challenge.
3. Notations and Task Formulation
We first introduce the notations used in this paper. Then, we describe the task definition of generative recommendation.
3.1. Notations
[…]
3.2. Generative Recommendation
The task of generative recommendation is given the input describing the information of $\mathcal{I}_{u}^{+}$, generating a list of GIDs as the recommendation result. The GID is generated through an auto-regressive manner. The probability of recommending item $i$ for user $u$ is estimated as:
(1) $p(u,i)=\prod_{t=1}^{l}p(z_{i}^{t}|z_{i}^{1},z_{i}^{2},\cdots,z_{i}^{t-1},\mathcal{I}_{u}^{+}).$
The recommender selects items with the top-$n$ highest $p(u,i)$ as the recommendation list for user $u$.
4. Methodology
In this section, we describe the details of ColaRec. We first provide overview of the proposed ColaRec. Then the construction of items’ GIDs are detailed. After that, the user-item recommendation task and the item-item indexing task are described. Finally, we describe the joint optimization of above tasks.
4.1. Overview of ColaRec
Figure 3 illustrates an overview of the proposed ColaRec. ColaRec constructs the GID using a graph-based CF model, which effectively captures collaborative signals. The training of ColaRec consists of two tasks: the user-item recommendation task and the item-item indexing task. The user-item recommendation aims to map the content information of the user’s historical interacted items into the GID of the recommended item. The item-item indexing task targets on the mapping from the item side information into the item’s GID. Both the two tasks are achieved through a shared encoder-decoder based language model to better capture the textual content information. To this end, the recommendation task unifies both collaborative signals and item content information for better recommendation, while the indexing task performs the alignment between collaborative signals and content information. Note that parameters of the involved language model are also fine-tuned to better adapt the language model for recommendation.
4.2. Generative Identifier Construction
The construction of GIDs plays a crucial role for generative recommendation. Generally speaking, GIDs should satisfy the following expectations for better recommendation: (i) GIDs need to contain knowledge about both collaborative signals and content information; (ii) correlated items (e.g., similar items in content or items interacted by similar users) should have correlated GIDs; (iii) each item should have one unique GID and each GID should correspond to one specific item.
To fulfil above expectations, we utilize a hierarchical clustering approach to construct GIDs from a graph-based CF model. Specifically, we first extract item representations from a pretrained LightGCN model. Then the constrained $K$-means algorithm is called hierarchically based on the item representations. For the $t$-th level clustering with $t\in[1,l-1]$, the number of clusters is set to $K^{l-t}$. The next level clustering is conducted with items in the current cluster as the whole instance set. Regarding the last level of leaf node, we randomly allocate 1 to $K$ to the item. In this way, we establish a $K$-ary tree to organize the item set. Each item corresponds to a leaf node, while the path from the root to the leaf node is the GID of the item. Since LightGCN is trained on the user-item interaction graph, the GID can naturally encodes collaborative signals. Meanwhile, there is a codebook embedding matrix for every position of the GID, which will incorporate the content information in the item indexing task. We give detailed description in section 4.4. To this end, the GID together with the corresponding codebook embedding helps the recommender to model both collaborative signals and content information.
4.3. User-Item Recommendation
Model Inputs. The input sequence of user-item recommendation for user $u$ consists of unordered tuples with each tuple describing the content information of one interacted item of this user. For textual description of item $i$, we adopt the universal data format from (Li et al., 2023b). Specifically, textual description $c_{i}$ of item $i$ is formulated as an item “sentence” that comes from a flatted attribute dictionary consisting of key-value attribute pairs $(k,v)$, i.e., $[k_{1}{:}v_{1},k_{2}{:}v_{2},\dots]$. Besides, we also introduce the item atomic identifier $iad_{i}$ into the content information to further increase model fidelity. To this end, the content tuple of item $i$ is formulated as:
(2) $c_{i}=[iad_{i},k_{1}{:}v_{1},k_{2}{:}v_{2},\dots].$
The key idea of CF is that users’ preferences can be inferred from their interacted items. Therefore, for each user $u$, the input consists of the content aggregation of item tuples that $u$ has interacted with to reinforce the collaborative signals.
Since the training of ColaRec has two tasks, in the user-item recommendation task, we augment the input by adding a textual “user prompt” (denoted as $prompt_{u}$) at the beginning of the input to inform the model that the ongoing task is the recommendation task. Thus, the input for the user-item recommendation task is:
(3) $X_{u}=[prompt_{u},\{c_{i}|i\in\mathcal{I}_{u}^{+}\}].$
Language Models for Recommendation. We employ an encoder-decoder based language model to capture textual content information. Given the model input $X_{u}$, the language model encoder captures the semantic information of $X_{u}$ and returns the hidden state $\text{Encoder}(X_{u})$. After that, given the generated tokens $z^{<t}$ before the $t$-th generation step, the decoder generates the latent representation $\mathbf{d}_{t}\in\mathbb{R}^{m}$ for the $t$-th token of GID. $m$ is the dimension of the latent representation. This process can be formulated as:
(4) $\mathbf{d}_{t}=\text{Decoder}(\text{Encoder}(X_{u}),z^{<t})$
The generation probability at step $t$ is estimated by $\mathbf{d}_{t}$ and the codebook embedding matrix for position $t$, which is formulated as:
(5) $p(z^{t}|z^{<t},X_{u})=\operatorname{softmax}(\mathbf{d}_{t}\cdot\mathbf{E}_{t}^{\top}),$
where $\mathbf{E}_{t}$ is the $t$-th step codebook embedding matrix.
We adopt the cross-entropy loss for model optimization. Specifically, given a $(u,i)$ pair in the training set, the generative loss of recommendation is formulated as:
(6) $\mathcal{L}_{\text{rec}}=-\sum_{t=1}^{l}\log p(z^{t}_{i}|X_{u},z^{1}_{i},z^{2}_{i},\cdots,z^{t-1}_{i}).$
In this work, we use a pretrained T5 (Raffel et al., 2020) as the language model. The parameters of T5 are also fine-tuned through back propagation to better adapt the language model for recommendation.
4.4. Item-Item Indexing
To align collaborative signals and item content information, we introduce an item-item indexing task which conducts the mapping from the content-based semantic space into the interaction-based collaborative space.
Model Inputs. The input sequence for item indexing contains the textual information of the item. Besides, we also introduce information of users who have interacted with this item, to further encode collaborative signals. Similar to the recommendation task, we augment the input an “item prompt” (denoted as $prompt_{i}$) at the beginning of the input. Therefore, the input of the indexing task is formulated as:
(7) $X_{i}=[prompt_{i},c_{i},\{uad_{u}|u\in\mathcal{U}_{i}^{+}\}].$
Language Models for Indexing. The indexing task is conducted through the same language model and codebook embeddings as the recommendation task. The generation probability for the indexing task is formulated similarly with Eq. (4) and Eq. (5) except that the model input is $X_{i}$ instead of $X_{u}$. We adopt the cross-entropy loss for parameter tuning. The loss for item indexing is defined as:
(8) $\mathcal{L}_{\text{index}}=-\sum_{t=1}^{l}\log p(z^{t}_{i}|X_{i},z^{1}_{i},z^{2}_{i},\cdots,z^{t-1}_{i}).$
Contrastive Optimization. To conduct better alignment between collaborative signals and content information, a contrastive loss is further introduced. The idea is that items with similar GIDs should also be similar in the content-based semantic space. To this end, for the item $i$, we randomly sample an item $i_{+}$ which has overlapped sub-sequence in GIDs as the positive sample, and another randomly sampled item $i_{-}$ without overlapped GID tokens as the negative sample. The contrastive loss is defined as:
(9) $\mathcal{L}_{\text{c}}=-\ln\sigma(\textbf{h}(X_{i})\cdot(\textbf{h}(X_{i_{+}})-\textbf{h}(X_{i_{-}}))),$
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$), and $\sigma$ denotes the sigmoid function. Such a contrastive loss helps the encoder to learn better item input representations.
4.5. Joint Optimization
Besides the above tasks, we further introduce a ranking loss to enhance the ranking ability of ColaRec. For a $(u,i)$ pair in the training dataset, we sample one item that the user $u$ has not interacted with, and for which the GID has no overlap with the positive item $i$, as the negative sample $i_{-}$. The BPR loss (Rendle et al., 2012) is utilized to optimize the ranking, which is formulated as:
(10) $\mathcal{L}_{\text{BPR}}=-\ln\sigma(\mathbf{h}(X_{u})\cdot(\mathbf{h}(X_{i})-\mathbf{h}(X_{i_{-}}))).$
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$). The above loss pushes together the positive $(u,i)$ pair in the semantic space and pushes away the negative $(u,i_{-})$ pair, thus helping the encoder to capture the ranking knowledge between items.
Finally, ColaRec is trained with the above-described tasks jointly:
(11) $\mathcal{L}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{index}}+\mathcal{L}_{\text{bpr}}+\alpha\mathcal{L}_{\text{c}},$
where $\alpha$ denotes the weight for the contrastive loss.
During the inference, recommendation lists are generated via beam search. To avoid the recommender from generating invalid GIDs that cannot be mapped to concrete items in the candidate set, we employ the constrained beam search (Cao et al., 2021) to limit the generative range of the current token based on the prefix tokens.
Table 1. Statistics of three public datasets after preprocessing.
Datasets #Users #Items #Interactions
Beauty 22,363 12,101 198,502
Sports 35,598 18,357 296,337
Phone 27,879 10,429 194,439
Table 2. Performance comparison on three public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. * denotes the paired t-test with significance p-value ¡ 0.1.
CF-based Methods Generative Methods
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
Beauty Recall@5 0.0447 0.0649 0.0551 0.0650 0.0530 0.0524 0.0128 0.0451 0.0519 0.0667*
Recall@10 0.0653 0.0952 0.0831 0.0940 0.0776 0.0741 0.0228 0.0705 0.0799 0.0993*
Recall@20 0.0889 0.1314 0.1193 0.1320 0.1093 0.1016 0.0360 0.1018 0.1154 0.1371*
NDCG@5 0.0315 0.0450 0.0377 0.0452 0.0362 0.0378 0.0084 0.0305 0.0350 0.0449
NDCG@10 0.0383 0.0549 0.0469 0.0547 0.0443 0.0450 0.0117 0.0385 0.0443 0.0556*
NDCG@20 0.0445 0.0643 0.0563 0.0646 0.0526 0.0521 0.0151 0.0470 0.0534 0.0654*
Sports Recall@5 0.0206 0.0418 0.0355 0.0427 0.0314 0.0273 0.0117 0.0320 0.0374 0.0442*
Recall@10 0.0321 0.0623 0.0557 0.0631 0.0476 0.0403 0.0178 0.0497 0.0572 0.0660*
Recall@20 0.0471 0.0901 0.0836 0.0908 0.0713 0.0569 0.0284 0.0766 0.0881 0.0964*
NDCG@5 0.0140 0.0288 0.0240 0.0294 0.0208 0.0193 0.0079 0.0225 0.0249 0.0294
NDCG@10 0.0177 0.0355 0.0306 0.0361 0.0261 0.0235 0.0099 0.0284 0.0313 0.0364
NDCG@20 0.0215 0.0426 0.0377 0.0431 0.0321 0.0278 0.0126 0.0350 0.0392 0.0442*
Phone Recall@5 0.0410 0.0713 0.0643 0.0717 0.0569 0.0470 0.0187 0.0412 0.0601 0.0745*
Recall@10 0.0603 0.1052 0.0976 0.1043 0.0855 0.0668 0.0341 0.0625 0.0895 0.1121*
Recall@20 0.0871 0.1487 0.1420 0.1481 0.1233 0.0928 0.0564 0.0966 0.1299 0.1587*
NDCG@5 0.0282 0.0481 0.0423 0.0488 0.0378 0.0315 0.0121 0.0282 0.0403 0.0490
NDCG@10 0.0344 0.0590 0.0530 0.0593 0.0470 0.0379 0.0170 0.0347 0.0498 0.0611*
NDCG@20 0.0412 0.0700 0.0643 0.0704 0.0566 0.0445 0.0225 0.0431 0.0600 0.0729*
5. Experiment Settings
In this section, we describe experiment settings to evaluate the proposed ColaRec. We aim to answer the following research questions:
  1. RQ1
    How does the proposed ColaRec perform compared with existing recommendation methods?
  2. RQ2
    How does the joint training of multiple tasks affect the performance of ColaRec?
  3. RQ3
    How does the design of GIDs affect the recommendation performance?
5.1. Datasets
We use three real-world public datasets from Amazon Product Reviews [^3] https://jmcauley.ucsd.edu/data/amazon/, one of the most widely used recommendation benchmarks, to evaluate the performance of ColaRec. In particular, the experiments are conducted on three subcategories, including “Beauty”, “Sports and Outdoors”, and “Cell Phones and Accessories”. Users and items which have less than five interactions are filtered out. Table 1 shows the statistics of all three datasets after preprocessing. As for content information, we select “title”, “brand” and “categories” from the item metadata as the textual content description of items.
5.2. Evaluation Protocols
We adopt cross-validation to evaluate the performance of recommnders. In this paper we target on general recommendation other than sequential recommendation. To this end, we randomly split each user’s historical interactions into the training/validation/test set with the ratio of 8:1:1. We employ two widely used metrics, recall@$n$ and normalized discount cumulative gain (NDCG@$n$), to evaluate the model performance. Recall evaluates how many ground-truth items occur in the recommended list, while NDCG further focuses on their rankings in the list. Note that in this paper, the candidate item set is the whole item set, other than a small subset with selected items. Each experiment is conducted three times and the average score is reported.
5.3. Baselines
We compare ColaRec with several representative related baselines, including both conventional CF-based methods and generative methods. CF-based baselines include NeuMF (He et al., 2017), LightGCN (He et al., 2020), SimpleX (Mao et al., 2021), and NCL (Lin et al., 2022). Generative baselines include MultiVAE (Liang et al., 2018), DiffRec (Wang et al., 2023b), DSI (Tay et al., 2022), and TIGER (Rajput et al., 2023).
Table 3. Performance comparison of long-tail users on three public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. ** denotes that the improvements are significant with p-value ¡ 0.05.
CF-based Methods Generative Methods
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
Beauty Recall@5 0.0416 0.0636 0.0555 0.0639 0.0510 0.0464 0.0131 0.0415 0.0519 0.0660**
Recall@10 0.0604 0.0922 0.0825 0.0907 0.0742 0.0662 0.0228 0.0653 0.0799 0.0975**
Recall@20 0.0817 0.1253 0.1160 0.1264 0.1039 0.0917 0.0354 0.0940 0.1154 0.1327**
Sports Recall@5 0.0209 0.0433 0.0355 0.0440 0.0329 0.0267 0.0116 0.0307 0.0380 0.0456**
Recall@10 0.0317 0.0639 0.0562 0.0645 0.0495 0.0394 0.0170 0.0472 0.0581 0.0674**
Recall@20 0.0468 0.0904 0.0836 0.0908 0.0725 0.0553 0.0273 0.0728 0.0882 0.0976**
Phone Recall@5 0.0405 0.0723 0.0660 0.0727 0.0571 0.0451 0.0206 0.0404 0.0602 0.0756**
Recall@10 0.0590 0.1054 0.0986 0.1043 0.0861 0.0641 0.0371 0.0623 0.0898 0.1131**
Recall@20 0.0855 0.1482 0.1418 0.1473 0.1228 0.0899 0.0600 0.0939 0.1293 0.1590**
  • NeuMF (He et al., 2017) enhances MF with multiple hidden layers to learn non-linear patterns in user-item interactions.
  • LightGCN (He et al., 2020) simplifies GNN for CF and learns user and item representations via linear neighborhood aggregation.
  • SimpleX (Mao et al., 2021) is a simple CF model with a cosine-based contrastive loss and negative sampling.
  • NCL (Lin et al., 2022) improves LightGCN with contrastive learning.
  • MultiVAE (Liang et al., 2018) is an autoencoder-based method, which utilizes VAEs to model the interaction signals.
  • DiffRec (Wang et al., 2023b) is a new proposed recommendation model based on diffusion models. DiffRec learns the user-item interaction knowledge through a recontruction and denoising manner.
  • DSI (Tay et al., 2022) is a generative document retrieval method. To adapt DSI for recommendation, we formulate the input as GIDs of the user’s historical interacted items. We use two versions of DSI. DSI-R refers to the DSI model with a random string as the GID of the item. DSI-S is a DSI model which constructs the item GID with a hierarchical $K$-means algorithm based on the item textual content embeddings from a pre-trained BERT model.
  • TIGER (Rajput et al., 2023) is a generative recommendation method. Specifically, a pre-trained Sentence-T5 encoder is used to obtain embeddings of the item’s textual content. These embeddings are then quantized using an RQ-VAE to build GIDs. We do not introduce sequential orders to adapt TIGER for general recommendation.
We don’t introduce P5-based baselines  (Geng et al., 2022; Hua et al., 2023) since these methods require the model input prompt to include candidate items for the general recommendation task. Given the limited input length, these methods cannot perform item ranking among the whole item set.
5.4. Implementation Details
[…]
Table 4. Ablation study on three datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
Beauty Sports Phone
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
(1) w/o textual content 0.0527 0.0809 0.0346 0.0439 0.0364 0.0542 0.0239 0.0297 0.0636 0.0974 0.0426 0.0535
(2) w/o indexing 0.0637 0.0947 0.0428 0.0531 0.0422 0.0644 0.0278 0.0350 0.0728 0.1086 0.0487 0.0602
(3) w/o $\mathcal{L}_{bpr}$ 0.0612 0.0918 0.0412 0.0512 0.0424 0.0634 0.0282 0.0350 0.0719 0.1077 0.0486 0.0602
(4) w/o $\mathcal{L}_{c}$ 0.0657 0.0989 0.0434 0.0544 0.0422 0.0649 0.0279 0.0352 0.0731 0.1106 0.0485 0.0606
Table 5. Comparison between different types of GIDs. The best and the second-best scores are marked in bold and underlined fonts, respectively. R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
Beauty Sports Phone
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
(1) $iad$ 0.0658 0.0983 0.0437 0.0544 0.0428 0.0657 0.0285 0.0359 0.0719 0.1074 0.0474 0.0589
(2) $Random$ 0.0600 0.0902 0.0401 0.0500 0.0411 0.0623 0.0272 0.0340 0.0667 0.1004 0.0443 0.0551
(3) $Content$ 0.0662 0.1008 0.0440 0.0554 0.0423 0.0647 0.0278 0.0350 0.0716 0.1080 0.0477 0.0593
5.5. Performance Comparison (RQ1)
To answer RQ1, we conduct a comparative analysis of the proposed ColaRec on both overall users and long-tail users.
5.5.1. Comparison on whole users
Table 2 shows the performance comparison on overall users. From these results, we make the following observations. Firstly, the proposed ColaRec achieves the best recommendation performance on all datasets except for NDCG@5 in Beauty, which achieves comparable scores with the best NCL baseline. In particular, ColaRec consistently outperforms previous CF-based and generative baselines in Recall@20, achieving a relative improvement of 3.87%, 6.17% and 6.72% on Beauty, Sports, and Phone, respectively. These results demonstrate the effectiveness of ColaRec and its generalization across different domains.
Secondly, compared with DSI-R and DSI-S, which directly adapt generative retrieval methods to the recommendation task, ColaRec achieves a notable 47.89%, 38.13%, and 44.70% relative improvement in terms of Recall@5 on three datasets respectively, demonstrating effectiveness of the proposed ColaRec. Such results also demonstrate that naively transferring the generative retrieval methods for recommendation cannot achieve satisfying results. Furthermore, our method consistently outperforms the state-of-the-art generative recommendation model TIGER, which uses a RQ-VAE module to construct GIDs. The reason may be that TIGER only considers item content information but the collaborative signals are overlooked without an explicit alignment process.
Lastly, existing generative methods (e.g., DiffRec and TIGER), while exhibiting promising results in some specific scenarios (Rajput et al., 2023; Wang et al., 2023b; Li et al., 2023a), still underperform the strong CF-based methods (e.g., NCL) in the general recommendation task. In contrast, ColaRec achieves competitive results compared to these CF methods on all datasets, demonstrating the potential of infusing content information for collaborative generation of recommender systems.
5.5.2. Comparison on long-tail users
We also conduct experiments to verify the recommendation performance of ColaRec on long-tail users with sparse interactions. In this experiment, the ratio between head users and long-tail users is set as 20%:80%. Table 3 reports the results of Recall on three datasets. Results of NDCG show similar trends and are omitted due to the reason of space. We can see that ColaRec significantly outperforms all baselines when generating recommendation for long-tail users. The reason is that ColaRec models both user-item interactions and item content information. Given the long-tail users with less interaction knowledge, the content information helps ColaRec to gain better performance.
To conclude, the proposed ColaRec is effective to yield better performance compared with existing baselines. This improvement is more significant on long-tail users.
5.6. Ablation Study (RQ2)
In this section, we conduct ablation studies to analyse the effectiveness of each component in ColaRec. We implement four ablative variants of ColaRec, including: (1) w/o textual contentdeletes all textual content information in the model input; (2) w/o indexingremoves the item-item indexing task; (3) w/o $\mathcal{L}_{bpr}$removes the ranking loss $\mathcal{L}_{\text{bpr}}$; and (4) w/o $\mathcal{L}_{c}$removes the contrastive loss $\mathcal{L}_{\text{c}}$.
The results on three datasets are shown in Table 4. These results clearly indicate that the removal of any component from our proposed method results in a noticeable decline in overall performance. From the results of variant (1), we see that removing textual content information leads to a significant reduction in performance. Specifically, a $26.57\%$, $21.43\%$, and $17.14\%$ drop in Recall@5 is observed across the three datasets. This suggests the importance of incorporating textual content information to enhance the models’ understanding of items. Furthermore, the variant (3), which is trained without $\mathcal{L}_{\text{bpr}}$, exhibits a notable performance decrease in comparison to ColaRec. This highlights the effectiveness of the pairwise ranking objective, which focuses on the relative prioritization of positive and negative items within the generative recommendation paradigm. Besides, the results of variant (2) and (4) verify the effectiveness of the proposed item-item indexing task and the contrastive objective $\mathcal{L}_{c}$. A consistent performance reduction is observed on the three datasets when either of the two techniques is removed. This suggests that aligning item content information with user-item collaborative signals not only facilitates mutual reinforcement but also enables the learning of more comprehensive and effective representations.
In conclusion, each component of ColaRec is essential to improve the recommendation performance.
(a) Beauty
(b) Sports
(c) Phone
Figure 4. Impact of the length of GIDs. The GID length $l$ is set from $1$ to $4$.
(a) Beauty
(b) Sports
(c) Phone
Figure 5. Impact of the number of clusters. The number of clusters $K$ varies from $32$ to $128$.
5.7. GID Investigation (RQ3)
The construction of GIDs plays a crucial role in generative recommendation as it defines the model’s search space for generation. To answer RQ3, we conduct two analytical experiments about GIDs: (1) an analysis of different GID types (e.g., the single token $iad$, the random assigned GID, and the content-based GID), and (2) an analysis of hyper-parameters when constructing GIDs (e.g., the length and $K$-means cluster numbers).
5.7.1. Effect of different GID types
To further evaluate our GID generation strategy, we conducted an ablation study comparing it with three techniques: iad-based GID, $Random$ GID, and $Content$ GID. Specifically, the iad-based GID assigns a unique single token to each item to represent it using the corresponding vector in the item embedding matrix. The Random GID assigns a random string to each item as the identifier, without considering any prior knowledge. Content GID represents the GID constructed from item content information. Unlike the collaborative GID used in ColaRec, the content-based GID employs a hierarchical $K$-means clustering algorithm (Tay et al., 2022) to cluster items based on their textual representation derived from a pretrained BERT. To make a fair comparison with ColaRec, the length and the codebook size of both $Random$ GIDs and $Content$ GIDs are identical to those in ColaRec.
Table 5 details our results on the three datasets, where ColaRec attains the best performance. This indicates that our GID construction method, based on collaborative signals, is highly effective for generative recommendation. The improvement over the Content GID highlights the effectiveness of collaborative signals in recommendation tasks. Furthermore, ColaRec’s superior performance compared to the iad method demonstrates the benefit of explicitly introducing item correlations into item GIDs. In addition, the Random method leads to the lowest performance as the random string could further introduce noisy signals in the learning process. These results illustrate the importance of constructing effective GIDs for generative recommenders.
5.7.2. Impact of GID Hyper-parameters
In this section, we investigate the impact of hyperparameters $l$ and $K$ in the GID construction process. Figure 4 shows the results of ColaRec with different GID lengths $l$, ranging from $1$ to $4$. We can see that as $l$ varies, the recommedation performance exhibits some fluctuations, with the best results typically observed when $l=3$. It is worth noting that the GID of ColaRec consists of two parts: (i) the prefix tokens $[z_{i}^{1},\dots,z_{i}^{l-1}]$ from hierarchy clustering; and (ii) the last token $z_{l}$, which is randomly assigned to ensure uniqueness of the GID. A longer GID tends to include a greater proportion of hierarchy clustering parts, potentially facilitating model learning. However, a too long GID means more autoregressive decoding steps in model generation, which increases search difficulty and inference latency. Therefore, balancing efficiency and effectiveness, we have selected $l=3$ as the default setting.
To assess how the number of clusters $K$ affects performance, we fixed the GID length $l$ as $l=3$ and varied $K$ with values $32$, $64$, $96$, and 128. As shown in Figure 5, a higher $K$ typically results in a slight decrease in overall performance. The decrease on Beauty is more notable. The reason is that a higher $K$ indicates a larger search space in the decoding process, and thus increase the generation difficulty. To this end, choosing a suitable clustering number according to number of items is important for generative recommendation.
6. Conclusions and future work
This paper proposes ColaRec, a novel framework to conduct content-based collaborative generation for recommender systems. As an end-to-end generative recommender, ColaRec effectively integrates both item content information and user-item collaborative signals within a unified framework through fine-tuning an encoder-decoder based language model. In addition, an auxiliary item indexing task and a contrastive loss are proposed to further align the model’s representations of item content and user-item collaborative signals. We have conducted extensive experiments on three real-world datasets and empirical results demonstrate the effectiveness of ColaRec.
In the future, we intend to investigate more methods to construct appropriate generative identifiers (GIDs) and adopt more effective approaches to better align content information and collaborative signals. We also plan to introduce larger language models together with larger volume of training data to generate better recommendation. Negative sampling for generative recommendation is one of our future works, either using GIDs to sample more informative negative samples or utilizing generative models to generate simulated negative instances. Besides, parameter efficient fine-tuning and model efficiency are also potential research directions.

### v2

@@
-Enhanced Generative Recommendation via Content and Collaboration Integration
+Content-Based Collaborative Generation for Recommender Systems
 Abstract.
-Generative recommendation has emerged as a promising paradigm aimed at augmenting recommender systems with recent advancements in generative artificial intelligence.
-This task has been formulated as a sequence-to-sequence generation process, wherein the input sequence encompasses data pertaining to the user’s previously interacted items, and the output sequence denotes the generative identifier for the suggested item.
-However, existing generative recommendation approaches still encounter challenges in (i) effectively integrating user-item collaborative signals and item content information within a unified generative framework, and (ii) executing an efficient alignment between content information and collaborative signals.
-In this paper, we introduce content-based colla-borative generation for recommender systems, denoted as ColaRec.
-To capture collaborative signals, the generative item identifiers are derived from a pretrained collaborative filtering model, while the user is represented through the aggregation of interacted items’ content.
-Subsequently, the aggregated textual description of items is fed into a language model to encapsulate content information.
-This integration enables ColaRec to amalgamate collaborative signals and content information within an end-to-end framework.
-Regarding the alignment, we propose an item indexing task to facilitate the mapping between the content-based semantic space and the interaction-based collaborative space.
-Additionally, a contrastive loss is introduced to ensure that items with similar collaborative GIDs possess comparable content representations, thereby enhancing alignment.
-To validate the efficacy of ColaRec, we conduct experiments on three benchmark datasets.
-Empirical results substantiate the superior performance of ColaRec.
-Recommender System, Generative Recommendation, Collaborative Filtering, Generative Model, Language Model
-isbn: ccs: Information systems Recommender systemsccs: Information systems Retrieval models and rankingccs: Information systems Novelty in information retrieval
+Generative models have emerged as a promising utility to enhance recommender systems.
+It is essential to model both item content and user-item collaborative interactions in a unified generative framework for better recommendation.
+Although some existing large language model (LLM)-based methods contribute to fusing content information and collaborative signals, they fundamentally rely on textual language generation, which is not fully aligned with the recommendation task.
+How to integrate content knowledge and collaborative interaction signals in a generative framework tailored for item recommendation is still an open research challenge.
+In this paper, we propose content-based colla-borative generation for recommender systems, namely ColaRec.
+ColaRec is a sequence-to-sequence framework which is tailored for directly generating the recommended item identifier.
+Precisely, the input sequence comprises data pertaining to the user’s interacted items, and the output sequence represents the generative identifier (GID) for the suggested item.
+To model collaborative signals, the GIDs are constructed from a pretrained collaborative filtering model, and the user is represented as the content aggregation of interacted items.
+To this end, ColaRec captures both collaborative signals and content information in a unified framework.
+Then an item indexing task is proposed to conduct the alignment between the content-based semantic space and the interaction-based collaborative space.
+Besides, a contrastive loss is further introduced to ensure that items with similar collaborative GIDs have similar content representations.
+To verify the effectiveness of ColaRec, we conduct experiments on four benchmark datasets.
+Empirical results demonstrate the superior performance of ColaRec.
+Recommender System, Generative Recommendation
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679692isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
 1. introduction
-Recommender systems are widely deployed to discover user interests and provide personalized information services (Hu et al., 2018; Qi et al., 2020; Yuan et al., 2019).
-Recently, generative models, such as large language models (LLMs) (Zhao et al., 2023; Chang et al., 2023) and diffusion models (Yang et al., 2023; Cao et al., 2022), have gained prominence in advancing artificial intelligence.
-In such a context, generative recommendation (Wang et al., 2023b; Bao et al., 2023; Si et al., 2023; Rajput et al., 2023; Geng et al., 2022; Ji et al., 2023) has emerged as a new paradigm to enhance recommendation agents with end-to-end generation capability of these models.
-Generative Recommendation.
-Generative recommendation assigns each item with a unique sequence of tokens as the item’s generative identifier, a.k.a., GID, then the recommendation is formulated as a sequence-to-sequence generation task, where the input is historical user-item interactions while the output sequence refers to the GID of the recommended item (Rajput et al., 2023; Si et al., 2023; Hua et al., 2023) [^1] Wang et al.
-(2023a) proposed another paradigm to directly generate new content, e.g., images, for recommendation.
-However, their methods are tailored for the generation of virtual content and cannot be used for recommendation of concrete items, such as the item recommendation in an e-commerce platform.
-In this paper, we target on the recommendation of concrete items..
-Figure 1. Comparison between conventional itemIDs and GIDs. GIDs contain more concrete correlations.
-Compared with conventional itemIDs with an assigned single random token, the sequential tokens of a GID contain more explicit representation information.
-As shown in Figure 1, we see correlated GIDs denote correlated items, and thus help the recommender to gain better performance.
-Generative recommendation provides an end-to-end paradigm to effectively utilize generative models for recommendation.
-Tay et al.
-(2022) initially introduced Differentiable Search Indexing (DSI), which marks an early endeavor in employing generative models for information retrieval.
-DSI has inspired a series of studies on generative recommendation (Rajput et al., 2023; Si et al., 2023; Petrov and Macdonald, 2023).
-The promising results of these works demonstrate the potential of generative recommendation.
-Limitations of Current Approaches. Existing generative recommendation approaches still have the following two limitations, leading to suboptimal performance.
-Firstly, existing methods fail to effectively model collaborative signals and item content information in a unified framework.
-Collaborative signals refer to the knowledge contained in the user-item interactions while item content information refers to the textual description of items, as shown in Figure 2.
-On one hand, several methods utilize either the item title or hierarchical content embeddings of items to construct GIDs (Li et al., 2023c; Rajput et al., 2023).
-However, these methods only consider the item content information while the collaborative signals between users and items are overlooked.
-On the other hand, Si et al.
-(2023) construct GIDs using item embeddings of a pretrained SASRec (Kang and McAuley, 2018).
-Although the GID in this work contains the item-item sequential (collaborative) connections, the proposed recommendation framework in their research fails to model the item content information.
-To achieve satisfying recommendation results, we argue that the generative recommenders should be able to jointly model the item content information and the user-item collaborative signals in a unified generative framework.
-Secondly, existing generative recommendation methods cannot effectively align item content information and collaborative signals.
-Although Hua et al.
-(2023) proposed to construct the GID using the concatenation of a semantic string learnt from the item content and a sequence of tokens learnt from the item-item co-occurrence matrix, the naive concatenation without a learning process is not effective to perform the alignment, leading to sub-optimal performance.
-The rich content information can assist the recommendation models in obtaining more informative and comprehensive item representations, thus enhancing the modeling of user-item collaborative interactions, and vice visa.
-To this end, the alignment between collaborative signals and content information is fundamental for further enhancing generative recommenders.
-However, we argue that a successful alignment should be achieved through an explicit learning process, e.g.,the mapping between the content-based semantic space and the interaction-based collaborative space.
 Figure 1.
@@
 Collaborative signals refer to the knowledge contained in user-item interactions while content information refers to the textual description of items.
-Proposed Methods.
-In this paper, we propose content-based colla-borative generation for recommender systems (ColaRec), a generative recommendation model which unifies both item content information and user-item collaborative interaction signals in a sequence-to-sequence generation framework.
-Taking a particular user as an example, the input sequence of ColaRec consists of unordered [^2] In this paper, we target on the general recommendation task other than sequential recommendation.
+Recommender systems are widely deployed to discover user interests and provide personalized information services (Hu et al., 2018; Qi et al., 2020; Yuan et al., 2019).
+Recently, generative models, such as large language models (LLMs) (Zhao et al., 2023; Chang et al., 2023) and diffusion models (Yang et al., 2023; Cao et al., 2022), have gained prominence in advancing artificial intelligence.
+In such a context, plenty of research has emerged to utilize generative models for recommendation (Wang et al., 2023b; Rajput et al., 2023).
+To achieve satisfying results, the recommender should be able to incorporate both user-item collaborative signals and item content information.
+Collaborative signals refer to the knowledge contained in the user-item interactions while item content information refers to the textual description of items, as shown in Figure 1.
+LLM-based recommendation is a straightforward solution to fuse content information with collaborative signals.
+In general, these approaches reorganize the historical user-item interactions into text-based natural language, aka the instruction-tuning dataset.
+Then, the recommendation task is reformulated as language generation under designed prompts  (Fan et al., 2023; Li et al., 2024).
+Although this kind of approach utilizes the powerful text generation capability of LLMs, there exists the inherent misalignment between the two tasks of language generation and item recommendation.
+For example, there is usually a non-trivial grounding stage to map the generated language to a concrete item (Lin et al., 2023; Bao et al., 2023).
+LLMs also suffer poor ranking performance to generate target itemIDs from a large candidate pool (Zhang et al., 2023b; Liao et al., 2023).
+The task discrepancy limits the practical usage of LLM-based methods.
+Inspired by generative retrieval (Tay et al., 2022), some works (Rajput et al., 2023; Si et al., 2023) assign each item with a unique sequence of tokens as the item’s generative identifier, aka GID, then a generative sequence-to-sequence model which is tailored for directly generating item recommendation is trained without the need of explicit language modeling.
+The input sequence comprises historical user-item interactions while the output sequence refers to the GID of the recommended item.
+In this paper, we refer such the above paradigm as Generative Recommendation [^1] Wang et al.
+(2023a) proposed another paradigm to directly generate new content, e.g., images, for recommendation.
+However, their methods are tailored for the generation of virtual content and cannot be used for concrete items.
+In this paper, we target on the recommendation of concrete items..
+Compared with conventional itemIDs with an assigned single random token, the sequential tokens of a GID contain more explicit representation information.
+As shown in Figure 2, we see correlated GIDs denote correlated items, and thus help the recommender to conduct more effective generation.
+Generative recommendation provides an end-to-end paradigm to design generative models tailored for the recommendation task.
+Plenty of studies have been conducted along this research line  (Rajput et al., 2023; Si et al., 2023; Petrov and Macdonald, 2023; Tan et al., 2024; Zheng et al., 2023).
+However, existing generative recommendation methods still fail to effectively model collaborative signals and content information in a unified framework.
+For example, several methods utilize either the item titles or hierarchical content embeddings of items to construct GIDs (Li et al., 2023c; Rajput et al., 2023).
+These methods only consider the item content information while the collaborative signals between users and items are overlooked.
+On the contrary, Si et al.
+(2023) constructed GIDs using item embeddings of a pretrained SASRec (Kang and McAuley, 2018).
+Although the GID in this work contains the item-item sequential (collaborative) connections, the proposed recommendation framework in their research fails to model the item content information.
+Figure 2. Comparison between conventional itemIDs and GIDs. GIDs contain more concrete correlations.
+Besides, existing generative recommendation methods cannot effectively align item content information and collaborative signals.
+Although Hua et al.
+(2023) proposed a combined approach that concatenates content-based semantic strings with collaborative IDs derived from the item-item co-occurrence matrix.
+However, this naive concatenation, lacking a proper learning process, fails to achieve effective alignment, resulting in sub-optimal performance.
+The alignment should be achieved through an explicit learning process, e.g., the mapping between the content-based semantic space and the interaction-based collaborative space.
+In this paper, we propose content-based colla-borative generation for recommender systems (ColaRec), a model that unifies both item content information and user-item collaborative signals in a sequence-to-sequence generation framework tailored for the recommendation task.
+Taking a particular user as an example, the input sequence of ColaRec consists of unordered [^2] In this paper, we focus on the general recommendation task other than sequential recommendation.
 To this end, we use unordered item tuples as the input.
 tuples with each tuple describing the content information of one interacted item of this user.
-Then an encoder-decoder framework is used to generate the GID of the target item.
-To address the first limitation, we propose two tailored designs to jointly model user-item collaborative signals and item content information.
+Then an encoder-decoder Transformer (Vaswani et al., 2017) is used to generate the GID of the target item.
+More precisely, we propose two tailored designs to jointly model user-item collaborative signals and item content information.
 Firstly, we propose to construct GIDs using item representations obtained from a pretrained collaborative filtering model.
 In this paper, we use LightGCN (He et al., 2020) as the pretrained model.
-The LightGCN model is trained on the user-item interaction graph and thus the constructed GID can effectively encodes the user-item collaborative signals.
+The LightGCN model is trained on the user-item interaction graph and thus the constructed GID can effectively encode the user-item collaborative signals.
 Note that the LightGCN model can also be alternated with other models.
@@
 This design also keeps inline with the nature of collaborative filtering.
-The aggregation of content-based tuples is fed to an encoder-decoder based language model to effectively capture the textual content information.
-In this paper, we use a fine-tuned T5 (Raffel et al., 2020) as the involved language model.
-We leave the investigation of larger language models as one of our future works.
-To address the second limitation, we propose an auxiliary item indexing task which targets on mapping the item side information into the GID of this item through the same encoder-decoder model.
-More precisely, the item side information contains both the textual content information and a set of users who have interacted with this item.
-To this end, the indexing task maps both the item content information and user-item interaction signals into the constructed GID, achieving better alignment between item content information and user-item collaborative signals.
+The aggregation of content-based tuples is fed to the generative Transformer model to effectively capture the item content information.
+To conduct the alignment between content information and collaborative signals, we propose an auxiliary item indexing task which targets on mapping the item side information into the GID of this item through the same encoder-decoder model.
+Specifically, the item side information contains both the textual content information and a set of users who have interacted with this item.
+To this end, the indexing task maps both the item content information and user-item interaction signals into the constructed GID, achieving better alignment.
 Besides, we further propose a contrastive loss to ensure that items with similar collaborative GIDs are also similar in the content-based semantic space.
-Such the contrastive loss also helps to perform alignment between content information and collaborative signals.
-To demonstrate the effectiveness of the proposed ColaRec, we conduct extensive experiments on three public accessible benchmark datasets.
-Experimental results show that the proposed ColaRec outperforms related state-of-the-art baselines.
-Main Contributions. Our main contributions are as follows:
-  • We propose ColaRec, an end-to-end generative recommender which utilizes an encoder-decoder based language model to jointly model item content information and user-item collaborative signals for recommendation.
+To demonstrate the effectiveness of the proposed ColaRec, we conduct extensive experiments on four public datasets.
+Experimental results show that the proposed ColaRec outperforms related state-of-the-art baselines [^3] The code of this work is available at https://github.com/Junewang0614/ColaRec.
+Our main contributions are as follows:
+  • We propose ColaRec, a generative recommendation framework which utilizes an encoder-decoder model to jointly capture content information and collaborative signals for recommendation.
   • We propose an auxiliary item indexing task and a contrastive loss to perform better alignment between item content information and user-item collaborative signals to further enhance the performance of generative recommendation.
-  • We conduct extensive experiments on three datasets to demonstrate the effectiveness of the proposed ColaRec. Experimental results show superior recommendation performance of ColaRec.
+  • We conduct extensive experiments on four datasets to demonstrate the effectiveness of the proposed ColaRec. Experimental results show superior recommendation performance of ColaRec.
 2. related work
-In this section, we review related prior literature on collaborative filtering and generative models.
+In this section, we review related literature on collaborative filtering and generative models for recommendation.
 2.1. Collaborative Filtering
 Collaborative filtering (CF) is one of the most representative methods to build a recommendation agent.
-CF believes that a user can be represented as the aggregation of her/his interacted items, and vice visa.
+CF believes that a user can be represented as the aggregation of her/his interacted items, and vice versa.
 The keystone to conduct collaborative filtering is the user-item interaction matrix.
@@
 (2023b) proposed RecFormer to model long text sequences for recommendation.
-Different with existing CF methods, in this paper we focus on the new paradigm of generative recommendation, where the recommended item is generated through the decoder of a sequence-to-sequence model.
-Generative recommendation provides new potential to utilize advances in the field of generative artificial intelligence to enhance the recommendation task.
-2.2. Generative Models
-Generative models have become a hot research topic to generate new content, such as images and text, from the collected data.
+Different from existing CF methods, in this paper we focus on the new paradigm of generative recommendation, where the item is represented as a sequence of tokens, i.e., GID, and the recommendation is provided in a generative fashion.
+2.2. Generative Models for Recommendation
+Generative models have become a hot research topic to generate new content, such as images and text.
 Variational Autoencoders (VAEs) (Kingma and Welling, 2014; Higgins et al., 2016) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Creswell et al., 2018), are two kinds of representative generative models.
-VAEs aim to learn a probabilistic mapping from the observed data to a latent space, and then leverage variational inference to approximate the posterior distribution of latent variables.
-GANs introduce a min-max gaming approach with a generator attempting to generate simulated samples, and the discriminator distinguishing between real and generated data.
+Besides, diffusion models (Ho et al., 2020; Yang et al., 2023; Cao et al., 2022) have also shown promising results in content generation.
+These generative models have also been utilized for recommender systems  (Liang et al., 2018; Shenbin et al., 2020; Cai and Cai, 2022; He et al., 2018; Guo et al., 2022; Wang et al., 2022b, 2023b; Li et al., 2023a).
 Recently, Transformers (Vaswani et al., 2017) have shown promise in language generation, leading to notable LLMs like GPTs (Generative Pre-trained Transformers).
-Besides, diffusion models (Ho et al., 2020; Yang et al., 2023; Cao et al., 2022) have also shown promising results in content generation.
-Generative models for document retrieval.
-Generative retrieval has drawn increasing attention to conduct document retrieval with a sequence-to-sequence framework, where the input sequence is the query and the output sequence is the document identifier.
-The bloom of LLMs encourages generative retrieval to better understand semantic patterns stored in the retrieval process.
+LLM-based recommendation provides a straightforward solution to fuse user-item interaction signals with item textual content  (Liao et al., 2023; Zheng et al., 2023; Zhang et al., 2023a; Luo et al., 2024; Zhu et al., 2023).
+Typically, these approaches need to construct the recommendation training data with natural language, i.e., the instruction-tuning dataset, then the recommendation task is transformed into text generation under designed language prompts  (Zhang et al., 2023b; Lin et al., 2023; Fan et al., 2023; Li et al., 2024).
+For instance, Luo et al.
+(2024) proposed to inject collaborative signals into LLM-based recommendation through prompt augmentation.
+(Zhang et al., 2023a) and (Liao et al., 2023) proposed to fuse item embeddings into prompt embeddings for CTR prediction and recommendation.
+Lin et al.
+(2023) utilized LLMs to generate multiple aspects of the recommended items including titles and attributes.
+LC-Rec (Zheng et al., 2023) proposed to perform recommendation through conducting various language generation tasks.
+Despite the strong generation capabilities of LLMs, there is an inherent misalignment between language generation and item recommendation, which limits the practical usage of LLM-based methods.
+For example, it is usually necessary to go through a non-trivial grounding stage to map the generated language into specific items (Bao et al., 2023; Lin et al., 2023; Liao et al., 2023).
+Besides, it is also challenging for LLMs to directly generate recommended items from a large candidate pool (Liao et al., 2023; Zhang et al., 2023b; Luo et al., 2024).
+Inspired by generative retrieval, generative recommendation is a new paradigm to design generative models which is tailored for the recommendation task.
+Generative retrieval, which relies on the parametric memory of generative models to directly generate relevant document identifiers, has drawn increasing attention.
 Tay et al.
-(2022) firstly proposed a differentiable search index (DSI) for document retrieval, which assigns each document with a structured semantic meaningful identifier.
-DSI can be regarded as the early attempt to conduct generative retrieval.
-Plenty of work has been conducted following this research line.
-For example, strings including titles (Tang et al., 2023; Cao et al., 2021), n-grams (Bevilacqua et al., 2022; Chen et al., 2023) or URLs (Zhou et al., 2022) are explored to construct the document identifier.
-Wang et al.
-(2022a) proposed NCI which produces a hierarchical document identifier to get fine-grained semantic mapping, and a novel decoder to consider the prefixes of document identifiers.
-GenRet (Sun et al., 2023) learns to tokenize the document into a short discrete representation via a discrete auto-encoding approach, and utilizes progressive training and diverse clustering to assist the model training process.
+(2022) firstly proposed a differentiable search index (DSI) for generative retrieval.
+Plenty of methods have been conducted following this research line.
+Representative works include SEAL (Bevilacqua et al., 2022), NCI (Wang et al., 2022a), and GenRet (Sun et al., 2023).
+Similar to generative retrieval, generative recommendation constructs a sequence of tokens, aka generative identifier (GID), for each item, then a generative sequence-to-sequence model which is tailored for directly generating item recommendation is trained without the need of explicit language modeling.
+The input comprises historical user-item interactions and the output refers to the GID of the recommended item.
+Following this paradigm, TIGER (Rajput et al., 2023) is a representative generative recommendation method which uses an RQ-VAE (Zeghidour et al., 2022) to construct the GID and then uses a encoder-decoder based transformer to generate sequential recommendation.
+Si et al.
+(2023) proposed to construct the GID from a pretrained SASRec (Kang and McAuley, 2018) model for sequential recommendation.
+Hua et al.
+(2023) further investigated the effect of item identifier construction.
+Tan et al.
+(2024) utilized human language tokens to represent each item with a unique textual identifier.
+Despite the rise of generative recommendation, existing generative recommendation methods still suffer from the ineffective infusion of collaborative signals and item content information.
+How to jointly model and align collaborative signals and content information in an end-to-end generative framework tailored for item recommendation is still an open research challenge.
 Figure 3.
@@
 Besides, a ranking loss $\mathcal{L}_{\text{bpr}}$ and a contrastive loss $\mathcal{L}_{\text{c}}$ are also introduced.
-A tuned T5 is used to accomplish these tasks.
-Generative models for recommendation.
-Generative models have also been utilized for recommender systems.
-Early works focus on utilizing VAEs (Liang et al., 2018; Shenbin et al., 2020; Cai and Cai, 2022) or GANs (He et al., 2018; Guo et al., 2022; Wang et al., 2022b) for recommendation.
-Recently, diffusion models have also been utilized to perform item recommendation (Wang et al., 2023b; Li et al., 2023a).
-Besides, LLM-based recommendation has also been proposed (Bao et al., 2023; Zhang et al., 2023; Petrov and Macdonald, 2023).
-These works explore the ability of LLMs for recommendation through parameter efficient tuning (Bao et al., 2023) or instruction tuning (Zhang et al., 2023).
-Meanwhile, several works attempt to convert various recommendation tasks into a unified natural language generation task, and train recommenders with multi-task optimization, like P5 (Geng et al., 2022) and M6-Rec (Cui et al., 2022).
-Notably, TIGER (Rajput et al., 2023) is a representative generative recommendation method which uses an RQ-VAE (Zeghidour et al., 2022) to construct the GID and then uses encoder-decoder based transformers to generate sequential recommendation.
-Si et al.
-(2023) proposed to construct the GID from a pretrained SASRec (Kang and McAuley, 2018) model for sequential recommendation.
-Hua et al.
-(2023) further investigated the effect of item identifier construction for generative recommendation.
-Despite the arise of generative recommenders, existing methods still suffer from the ineffective infusion of collaborative signals and item content information.
-How to jointly model and align collaborative signals and content information in a sequence-to-sequence generative framework for recommender systems is still an open research challenge.
 3. Notations and Task Formulation
-We first introduce the notations used in this paper. Then, we describe the task definition of generative recommendation.
 3.1. Notations
@@
 The probability of recommending item $i$ for user $u$ is estimated as:
-(1) $p(u,i)=\prod_{t=1}^{l}p(z_{i}^{t}|z_{i}^{1},z_{i}^{2},\cdots,z_{i}^{t-1},\mathcal{I}_{u}^{+}).$
+(1) $p(u,i)=\prod_{t=1}^{l}p(z_{i}^{t}|\mathcal{I}_{u}^{+},z_{i}^{1},z_{i}^{2},\cdots,z_{i}^{t-1}).$
 The recommender selects items with the top-$n$ highest $p(u,i)$ as the recommendation list for user $u$.
@@
 In this section, we describe the details of ColaRec.
-We first provide overview of the proposed ColaRec.
-Then the construction of items’ GIDs are detailed.
+We first provide the overview of the proposed ColaRec.
+Then the construction of items’ GIDs is detailed.
 After that, the user-item recommendation task and the item-item indexing task are described.
-Finally, we describe the joint optimization of above tasks.
+Finally, we describe the joint optimization of the above tasks.
 4.1. Overview of ColaRec
@@
 The item-item indexing task targets on the mapping from the item side information into the item’s GID.
-Both the two tasks are achieved through a shared encoder-decoder based language model to better capture the textual content information.
-To this end, the recommendation task unifies both collaborative signals and item content information for better recommendation, while the indexing task performs the alignment between collaborative signals and content information.
-Note that parameters of the involved language model are also fine-tuned to better adapt the language model for recommendation.
+Both the two tasks are achieved through a shared encoder-decoder based model.
+The recommendation task unifies both collaborative signals and item content information for better recommendation, while the indexing task performs the alignment between collaborative signals and content information.
 4.2. Generative Identifier Construction
-The construction of GIDs plays a crucial role for generative recommendation.
+The construction of GIDs plays a crucial role in generative recommendation.
 Generally speaking, GIDs should satisfy the following expectations for better recommendation: (i) GIDs need to contain knowledge about both collaborative signals and content information; (ii) Correlated items (e.g., similar items in content or items interacted by similar users) should have correlated GIDs; (iii) Each item should have one unique GID and each GID should correspond to one specific item.
-To fulfil above expectations, we utilize a hierarchical clustering approach to construct GIDs from a graph-based CF model.
+To fulfill above expectations, we utilize a hierarchical clustering approach to construct GIDs from a graph-based CF model.
 Specifically, we first extract item representations from a pretrained LightGCN model.
 Then the constrained $K$-means algorithm is called hierarchically based on the item representations.
-For the $t$-th level clustering with $t\in[1,l-1]$, the number of clusters is set to $K^{l-t}$.
-The next level clustering is conducted with items in the current cluster as the whole instance set.
-Regarding the last level of leaf node, we randomly allocate 1 to $K$ to the item.
+The next level clustering is conducted with items in the current cluster as the whole item set.
+For the $t$-th level clustering with $t\in[1,l-1]$, the number of items is no more than $K^{l-t}$ in each cluster.
+Regarding the last level of leaf nodes, we randomly allocate 1 to $K$ to the items.
 In this way, we establish a $K$-ary tree to organize the item set.
 Each item corresponds to a leaf node, while the path from the root to the leaf node is the GID of the item.
-Since LightGCN is trained on the user-item interaction graph, the GID can naturally encodes collaborative signals.
+Since LightGCN is trained on the user-item interaction graph, the GID can naturally encode collaborative signals.
 Meanwhile, there is a codebook embedding matrix for every position of the GID, which will incorporate the content information in the item indexing task.
-We give detailed description in section 4.4.
-To this end, the GID together with the corresponding codebook embedding helps the recommender to model both collaborative signals and content information.
+We give the detailed description in section 4.4.
+To this end, the GID together with the corresponding codebook embeddings helps the recommender to model both collaborative signals and content information.
 4.3. User-Item Recommendation
@@
 The input sequence of user-item recommendation for user $u$ consists of unordered tuples with each tuple describing the content information of one interacted item of this user.
-For textual description of item $i$, we adopt the universal data format from (Li et al., 2023b).
-Specifically, textual description $c_{i}$ of item $i$ is formulated as an item “sentence” that comes from a flatted attribute dictionary consisting of key-value attribute pairs $(k,v)$, i.e., $[k_{1}{:}v_{1},k_{2}{:}v_{2},\dots]$.
+For the textual description of item $i$, we adopt the universal data format from (Li et al., 2023b).
+Specifically, textual description $c_{i}$ of item $i$ is formulated as a sequence that comes from a flatted attribute dictionary consisting of key-value attribute pairs $(k,v)$, i.e., $[k_{1}{:}v_{1},k_{2}{:}v_{2},\dots]$.
 Besides, we also introduce the item atomic identifier $iad_{i}$ into the content information to further increase model fidelity.
@@
 Therefore, for each user $u$, the input consists of the content aggregation of item tuples that $u$ has interacted with to reinforce the collaborative signals.
-Since the training of ColaRec has two tasks, in the user-item recommendation task, we augment the input by adding a textual “user prompt” (denoted as $prompt_{u}$) at the beginning of the input to inform the model that the ongoing task is the recommendation task.
+Since the training of ColaRec has two tasks, in the user-item recommendation task, we augment the input with a special task token $task_{u}$ at the beginning of the input to inform the model that the ongoing task is the recommendation task.
 Thus, the input for the user-item recommendation task is:
-(3) $X_{u}=[prompt_{u},\{c_{i}|i\in\mathcal{I}_{u}^{+}\}].$
-Language Models for Recommendation.
-We employ an encoder-decoder based language model to capture textual content information.
-Given the model input $X_{u}$, the language model encoder captures the semantic information of $X_{u}$ and returns the hidden state $\text{Encoder}(X_{u})$.
+(3) $X_{u}=[task_{u},\{c_{i}|i\in\mathcal{I}_{u}^{+}\}].$
+Item Generation.
+We employ an encoder-decoder based Transformer model to generate item recommendation.
+Given the model input $X_{u}$, the model encoder captures the semantic information of $X_{u}$ and returns the hidden state $\text{Encoder}(X_{u})$.
 After that, given the generated tokens $z^{<t}$ before the $t$-th generation step, the decoder generates the latent representation $\mathbf{d}_{t}\in\mathbb{R}^{m}$ for the $t$-th token of GID.
@@
 (6) $\mathcal{L}_{\text{rec}}=-\sum_{t=1}^{l}\log p(z^{t}_{i}|X_{u},z^{1}_{i},z^{2}_{i},\cdots,z^{t-1}_{i}).$
-In this work, we use a pretrained T5 (Raffel et al., 2020) as the language model.
-The parameters of T5 are also fine-tuned through back propagation to better adapt the language model for recommendation.
+In this work, we use a pretrained T5 (Raffel et al., 2020) as the Transformer model.
+The parameters of T5 are also fine-tuned through back propagation to better adapt the model for recommendation.
+Note that our approach is model-agnostic and T5 can be replaced by other sequence-to-sequence models, even without pretraining.
 4.4. Item-Item Indexing
@@
 Besides, we also introduce information of users who have interacted with this item, to further encode collaborative signals.
-Similar to the recommendation task, we augment the input an “item prompt” (denoted as $prompt_{i}$) at the beginning of the input.
-Therefore, the input of the indexing task is formulated as:
-(7) $X_{i}=[prompt_{i},c_{i},\{uad_{u}|u\in\mathcal{U}_{i}^{+}\}].$
-Language Models for Indexing.
-The indexing task is conducted through the same language model and codebook embeddings as the recommendation task.
+Similar to the recommendation task, we augment the input with a special task token $task_{i}$ at the beginning of the input for item indexing.
+Therefore, the input of item indexing is formulated as:
+(7) $X_{i}=[task_{i},c_{i},\{uad_{u}|u\in\mathcal{U}_{i}^{+}\}].$
+Item Indexing.
+The indexing task is conducted through the same model and codebook embeddings as the recommendation task.
 The generation probability for the indexing task is formulated similarly with Eq.
@@
 (5) except that the model input is $X_{i}$ instead of $X_{u}$.
-We adopt the cross-entropy loss for parameter tuning.
+We adopt the cross-entropy loss for parameter learning.
 The loss for item indexing is defined as:
 (8) $\mathcal{L}_{\text{index}}=-\sum_{t=1}^{l}\log p(z^{t}_{i}|X_{i},z^{1}_{i},z^{2}_{i},\cdots,z^{t-1}_{i}).$
-Contrastive Optimization.
+4.5. Multi-Task Training
+Besides the above two tasks, we further introduce a ranking loss to enhance the ranking ability of ColaRec and a contrastive loss to conduct better alignment.
+Item Ranking.
+For a $(u,i)$ pair in the training dataset, we randomly sample one item that the user $u$ has not interacted with as the negative sample $i_{-}$.
+The BPR loss (Rendle et al., 2012) is utilized to optimize the item ranking, which is formulated as:
+(9) $\mathcal{L}_{\text{bpr}}=-\ln\sigma(\mathbf{h}(X_{u})\cdot(\mathbf{h}(X_{i})-\mathbf{h}(X_{i_{-}}))),$
+where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$), and $\sigma$ denotes the sigmoid function.
+The above loss pushes together the positive $(u,i)$ pair and pushes away the negative $(u,i_{-})$ pair, thus helping the model to capture the ranking knowledge between items.
+Contrastive Learning.
 To conduct better alignment between collaborative signals and content information, a contrastive loss is further introduced.
 The idea is that items with similar GIDs should also be similar in the content-based semantic space.
-To this end, for the item $i$, we randomly sample an item $i_{+}$ which has overlapped sub-sequence in GIDs as the positive sample, and another randomly sampled item $i_{-}$ without overlapped GID tokens as the negative sample.
+To this end, for the item $i$, we randomly sample an item $i_{+}$ which has the overlapped prefix tokens in GIDs as the positive sample.
+The sampled item $i_{-}$ in $\mathcal{L}_{\text{bpr}}$ serves as the negative sample.
+Note that here we ensure the sampled item $i_{-}$ in $\mathcal{L}_{\text{bpr}}$ has no overlapped GID tokens with item $i$.
 The contrastive loss is defined as:
 (10) $\mathcal{L}_{\text{c}}=-\ln\sigma(\textbf{h}(X_{i})\cdot(\textbf{h}(X_{i_{+}})-\textbf{h}(X_{i_{-}}))),$
-where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$), and $\sigma$ denotes the sigmoid function.
-Such a contrastive loss helps the encoder to learn better item input representations.
-4.5. Joint Optimization
-Besides the above tasks, we further introduce a ranking loss to enhance the ranking ability of ColaRec.
-For a $(u,i)$ pair in the training dataset, we sample one item that the user $u$ has not interacted with, and for which the GID has no overlap with the positive item $i$, as the negative sample $i_{-}$.
-The BPR loss (Rendle et al., 2012) is utilized to optimize the ranking, which is formulated as:
-(10) $\mathcal{L}_{\text{BPR}}=-\ln\sigma(\mathbf{h}(X_{u})\cdot(\mathbf{h}(X_{i})-\mathbf{h}(X_{i_{-}}))).$
-where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$).
-The above loss pushes together the positive $(u,i)$ pair in the semantic space and pushes away the negative $(u,i_{-})$ pair, thus helping the encoder to capture the ranking knowledge between items.
-Finally, ColaRec is trained with the above-described tasks jointly:
+Such a contrastive loss helps the model to learn better item input representations.
+Joint Optimization. Finally, ColaRec is trained with the above described tasks jointly:
 (11) $\mathcal{L}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{index}}+\mathcal{L}_{\text{bpr}}+\alpha\mathcal{L}_{\text{c}},$
 where $\alpha$ denotes the weight for the contrastive loss.
-During the inference, recommendation lists are generated via beam search.
-To avoid the recommender from generating invalid GIDs that cannot be mapped to concrete items in the candidate set, we employ the constrained beam search (Cao et al., 2021) to limit the generative range of the current token based on the prefix tokens.
-Table 1. Statistics of three public datasets after preprocessing.
+During the inference, to avoid the recommender from generating invalid GIDs, we employ the constrained beam search (Cao et al., 2021) to limit the generative range of the current token based on the prefix tokens.
+Table 1. Statistics of four public datasets after preprocessing.
 Datasets #Users #Items #Interactions
@@
 Phone 27,879 10,429 194,439
+Recipe 17,813 41,240 555,618
 Table 2.
-Performance comparison on three public datasets.
+Performance comparison on four public datasets.
 The best and the second-best scores are marked in bold and underlined fonts, respectively.
 * denotes the paired t-test with significance p-value ¡ 0.1.
-CF-based Methods Generative Methods
-Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
-Beauty Recall@5 0.0447 0.0649 0.0551 0.0650 0.0530 0.0524 0.0128 0.0451 0.0519 0.0667*
-Recall@10 0.0653 0.0952 0.0831 0.0940 0.0776 0.0741 0.0228 0.0705 0.0799 0.0993*
-Recall@20 0.0889 0.1314 0.1193 0.1320 0.1093 0.1016 0.0360 0.1018 0.1154 0.1371*
-NDCG@5 0.0315 0.0450 0.0377 0.0452 0.0362 0.0378 0.0084 0.0305 0.0350 0.0449
-NDCG@10 0.0383 0.0549 0.0469 0.0547 0.0443 0.0450 0.0117 0.0385 0.0443 0.0556*
-NDCG@20 0.0445 0.0643 0.0563 0.0646 0.0526 0.0521 0.0151 0.0470 0.0534 0.0654*
-Sports Recall@5 0.0206 0.0418 0.0355 0.0427 0.0314 0.0273 0.0117 0.0320 0.0374 0.0442*
-Recall@10 0.0321 0.0623 0.0557 0.0631 0.0476 0.0403 0.0178 0.0497 0.0572 0.0660*
-Recall@20 0.0471 0.0901 0.0836 0.0908 0.0713 0.0569 0.0284 0.0766 0.0881 0.0964*
-NDCG@5 0.0140 0.0288 0.0240 0.0294 0.0208 0.0193 0.0079 0.0225 0.0249 0.0294
-NDCG@10 0.0177 0.0355 0.0306 0.0361 0.0261 0.0235 0.0099 0.0284 0.0313 0.0364
-NDCG@20 0.0215 0.0426 0.0377 0.0431 0.0321 0.0278 0.0126 0.0350 0.0392 0.0442*
-Phone Recall@5 0.0410 0.0713 0.0643 0.0717 0.0569 0.0470 0.0187 0.0412 0.0601 0.0745*
-Recall@10 0.0603 0.1052 0.0976 0.1043 0.0855 0.0668 0.0341 0.0625 0.0895 0.1121*
-Recall@20 0.0871 0.1487 0.1420 0.1481 0.1233 0.0928 0.0564 0.0966 0.1299 0.1587*
-NDCG@5 0.0282 0.0481 0.0423 0.0488 0.0378 0.0315 0.0121 0.0282 0.0403 0.0490
-NDCG@10 0.0344 0.0590 0.0530 0.0593 0.0470 0.0379 0.0170 0.0347 0.0498 0.0611*
-NDCG@20 0.0412 0.0700 0.0643 0.0704 0.0566 0.0445 0.0225 0.0431 0.0600 0.0729*
-5. Experiment Settings
-In this section, we describe experiment settings to evaluate the proposed ColaRec. We aim to answer the following research questions:
+R and N stand for Recall and NDCG.
+CF-based Methods Generative Models for Recommendation
+Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER LC-Rec Ours
+Beauty R@5 0.0447 0.0649 0.0551 0.0650 0.0530 0.0524 0.0128 0.0451 0.0519 0.0492 0.0667*
+R@10 0.0653 0.0952 0.0831 0.0940 0.0776 0.0741 0.0228 0.0705 0.0799 0.0770 0.0993*
+R@20 0.0889 0.1314 0.1193 0.1320 0.1093 0.1016 0.0360 0.1018 0.1154 0.1104 0.1371*
+N@5 0.0315 0.0450 0.0377 0.0452 0.0362 0.0378 0.0084 0.0305 0.0350 0.0326 0.0449
+N@10 0.0383 0.0549 0.0469 0.0547 0.0443 0.0450 0.0117 0.0385 0.0443 0.0415 0.0556*
+N@20 0.0445 0.0643 0.0563 0.0646 0.0526 0.0521 0.0151 0.0470 0.0534 0.0499 0.0654*
+Sports R@5 0.0206 0.0418 0.0355 0.0427 0.0314 0.0273 0.0117 0.0320 0.0374 0.0397 0.0442*
+R@10 0.0321 0.0623 0.0557 0.0631 0.0476 0.0403 0.0178 0.0497 0.0572 0.0617 0.0660*
+R@20 0.0471 0.0901 0.0836 0.0908 0.0713 0.0569 0.0284 0.0766 0.0881 0.0931 0.0964*
+N@5 0.0140 0.0288 0.0240 0.0294 0.0208 0.0193 0.0079 0.0225 0.0249 0.0264 0.0294
+N@10 0.0177 0.0355 0.0306 0.0359 0.0261 0.0235 0.0099 0.0284 0.0313 0.0335 0.0364*
+N@20 0.0215 0.0426 0.0377 0.0431 0.0321 0.0278 0.0126 0.0350 0.0392 0.0413 0.0442*
+Phone R@5 0.0410 0.0713 0.0643 0.0717 0.0569 0.0470 0.0187 0.0412 0.0601 0.0615 0.0745*
+R@10 0.0603 0.1052 0.0976 0.1043 0.0855 0.0668 0.0341 0.0625 0.0895 0.0919 0.1121*
+R@20 0.0871 0.1487 0.1420 0.1481 0.1233 0.0928 0.0564 0.0966 0.1299 0.1354 0.1587*
+N@5 0.0282 0.0481 0.0423 0.0486 0.0378 0.0315 0.0121 0.0282 0.0403 0.0408 0.0490*
+N@10 0.0344 0.0590 0.0530 0.0593 0.0470 0.0379 0.0170 0.0347 0.0498 0.0506 0.0611*
+N@20 0.0412 0.0700 0.0643 0.0704 0.0566 0.0445 0.0225 0.0431 0.0600 0.0615 0.0729*
+Recipe R@5 0.0118 0.0188 0.0114 0.0192 0.0167 0.0142 0.0142 0.0157 0.0168 0.0174 0.0198*
+R@10 0.0210 0.0296 0.0202 0.0298 0.0285 0.0235 0.0248 0.0270 0.0292 0.0289 0.0306*
+R@20 0.0339 0.0454 0.0328 0.0459 0.0462 0.0343 0.0403 0.0436 0.0464 0.0454 0.0482*
+N@5 0.0088 0.0149 0.0093 0.0149 0.0128 0.0105 0.0107 0.0122 0.0137 0.0138 0.0151*
+N@10 0.0119 0.0182 0.0122 0.0182 0.0167 0.0135 0.0141 0.0158 0.0176 0.0175 0.0185*
+N@20 0.0154 0.0223 0.0156 0.0224 0.0214 0.0165 0.0182 0.0202 0.0221 0.0218 0.0232*
+5. Experiment
+In this section, we conduct experiments to evaluate the proposed ColaRec. We aim to answer the following research questions:
   1. RQ1
@@
 5.1. Datasets
-We use three real-world public datasets from Amazon Product Reviews [^3] https://jmcauley.ucsd.edu/data/amazon/, one of the most widely used recommendation benchmarks, to evaluate the performance of ColaRec.
-In particular, the experiments are conducted on three subcategories, including “Beauty”, “Sports and Outdoors”, and “Cell Phones and Accessories”.
-Users and items which have less than five interactions are filtered out.
-Table 1 shows the statistics of all three datasets after preprocessing.
-As for content information, we select “title”, “brand” and “categories” from the item metadata as the textual content description of items.
+We use four real-world public datasets to evaluate the performance of ColaRec.
+Specifically, the experiments are conducted on three subcategories from Amazon Product Reviews [^4] https://jmcauley.ucsd.edu/data/amazon/ (“Beauty”, “Sports and Outdoors”, and “Cell Phones and Accessories”) and “Recipe” from Food.com [^5] https://www.kaggle.com/datasets/shuyangli94/food-com-recipes-and-user-interactions.
+Users and items that have less than five interactions are filtered out.
+Table 1 shows the statistics of all four datasets.
+As for content information, we use “title”, “brand” and “categories” from the Amazon item metadata as the textual content information of items.
+For Recipe, we use “name”, “description”, and “tag” to describe item content.
 5.2. Evaluation Protocols
-We adopt cross-validation to evaluate the performance of recommnders.
-In this paper we target on general recommendation other than sequential recommendation.
+We adopt cross-validation to evaluate the performance of recommenders.
+In this paper we focus on general recommendation rather than sequential recommendation.
 To this end, we randomly split each user’s historical interactions into the training/validation/test set with the ratio of 8:1:1.
@@
 5.3. Baselines
-We compare ColaRec with several representative related baselines, including both conventional CF-based methods and generative methods.
+We compare ColaRec with several representative related baselines, including both conventional CF-based methods and generative models for recommendation.
 CF-based baselines include NeuMF (He et al., 2017), LightGCN (He et al., 2020), SimpleX (Mao et al., 2021), and NCL (Lin et al., 2022).
-Generative baselines include MultiVAE (Liang et al., 2018), DiffRec (Wang et al., 2023b), DSI (Tay et al., 2022), and TIGER (Rajput et al., 2023).
-Table 3.
-Performance comparison of long-tail users on three public datasets.
-The best and the second-best scores are marked in bold and underlined fonts, respectively.
-** denotes that the improvements are significant with p-value ¡ 0.05.
-CF-based Methods Generative Methods
-Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
-Beauty Recall@5 0.0416 0.0636 0.0555 0.0639 0.0510 0.0464 0.0131 0.0415 0.0519 0.0660**
-Recall@10 0.0604 0.0922 0.0825 0.0907 0.0742 0.0662 0.0228 0.0653 0.0799 0.0975**
-Recall@20 0.0817 0.1253 0.1160 0.1264 0.1039 0.0917 0.0354 0.0940 0.1154 0.1327**
-Sports Recall@5 0.0209 0.0433 0.0355 0.0440 0.0329 0.0267 0.0116 0.0307 0.0380 0.0456**
-Recall@10 0.0317 0.0639 0.0562 0.0645 0.0495 0.0394 0.0170 0.0472 0.0581 0.0674**
-Recall@20 0.0468 0.0904 0.0836 0.0908 0.0725 0.0553 0.0273 0.0728 0.0882 0.0976**
-Phone Recall@5 0.0405 0.0723 0.0660 0.0727 0.0571 0.0451 0.0206 0.0404 0.0602 0.0756**
-Recall@10 0.0590 0.1054 0.0986 0.1043 0.0861 0.0641 0.0371 0.0623 0.0898 0.1131**
-Recall@20 0.0855 0.1482 0.1418 0.1473 0.1228 0.0899 0.0600 0.0939 0.1293 0.1590**
+Generative models for recommendation include MultiVAE (Liang et al., 2018), DiffRec (Wang et al., 2023b), DSI (Tay et al., 2022), TIGER (Rajput et al., 2023) and LC-Rec(Zheng et al., 2023).
   • NeuMF (He et al., 2017) enhances MF with multiple hidden layers to learn non-linear patterns in user-item interactions.
@@
   • MultiVAE (Liang et al., 2018) is an autoencoder-based method, which utilizes VAEs to model the interaction signals.
-  • DiffRec (Wang et al., 2023b) is a new proposed recommendation model based on diffusion models. DiffRec learns the user-item interaction knowledge through a recontruction and denoising manner.
+  • DiffRec (Wang et al., 2023b) is a new proposed recommendation model based on diffusion models. DiffRec learns the user-item interaction knowledge through a reconstruction and denoising manner.
   • DSI (Tay et al., 2022) is a generative document retrieval method.
@@
 DSI-R refers to the DSI model with a random string as the GID of the item.
-DSI-S is a DSI model which constructs the item GID with a hierarchical $K$-means algorithm based on the item textual content embeddings from a pre-trained BERT model.
+DSI-S is a DSI model which constructs the item GIDs with a hierarchical $K$-means algorithm based on the item textual content embeddings from a pretrained BERT model.
   • TIGER (Rajput et al., 2023) is a generative recommendation method.
-Specifically, a pre-trained Sentence-T5 encoder is used to obtain embeddings of the item’s textual content.
+Specifically, a pretrained Sentence-T5 encoder is used to obtain embeddings of the items’ textual content.
 These embeddings are then quantized using an RQ-VAE to build GIDs.
 We do not introduce sequential orders to adapt TIGER for general recommendation.
+  • LC-Rec (Zheng et al., 2023) follows a similar approach to TIGER (Rajput et al., 2023) for item index learning.
+Besides, it employs various language generation tasks under different prompts to accomplish recommendation.
 We don’t introduce P5-based baselines  (Geng et al., 2022; Hua et al., 2023) since these methods require the model input prompt to include candidate items for the general recommendation task.
@@
 […]
-Table 4.
-Ablation study on three datasets.
-The best and the second-best scores are marked in bold and underlined fonts, respectively.
-R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
-Beauty Sports Phone
-R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
-ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
-(1) w/o textual content 0.0527 0.0809 0.0346 0.0439 0.0364 0.0542 0.0239 0.0297 0.0636 0.0974 0.0426 0.0535
-(2) w/o indexing 0.0637 0.0947 0.0428 0.0531 0.0422 0.0644 0.0278 0.0350 0.0728 0.1086 0.0487 0.0602
-(3) w/o $\mathcal{L}_{bpr}$ 0.0612 0.0918 0.0412 0.0512 0.0424 0.0634 0.0282 0.0350 0.0719 0.1077 0.0486 0.0602
-(4) w/o $\mathcal{L}_{c}$ 0.0657 0.0989 0.0434 0.0544 0.0422 0.0649 0.0279 0.0352 0.0731 0.1106 0.0485 0.0606
-Table 5.
-Comparison between different types of GIDs.
-The best and the second-best scores are marked in bold and underlined fonts, respectively.
-R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
-Beauty Sports Phone
-R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
-ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
-(1) $iad$ 0.0658 0.0983 0.0437 0.0544 0.0428 0.0657 0.0285 0.0359 0.0719 0.1074 0.0474 0.0589
-(2) $Random$ 0.0600 0.0902 0.0401 0.0500 0.0411 0.0623 0.0272 0.0340 0.0667 0.1004 0.0443 0.0551
-(3) $Content$ 0.0662 0.1008 0.0440 0.0554 0.0423 0.0647 0.0278 0.0350 0.0716 0.1080 0.0477 0.0593
 5.5. Performance Comparison (RQ1)
@@
 5.5.1. Comparison on whole users
-Table 2 shows the performance comparison on overall users.
+Table 2 shows the performance comparison of overall users.
 From these results, we make the following observations.
-Firstly, the proposed ColaRec achieves the best recommendation performance on all datasets except for NDCG@5 in Beauty, which achieves comparable scores with the best NCL baseline.
-In particular, ColaRec consistently outperforms previous CF-based and generative baselines in Recall@20, achieving a relative improvement of 3.87%, 6.17% and 6.72% on Beauty, Sports, and Phone, respectively.
+Firstly, ColaRec achieves the best recommendation performance on all datasets except for NDCG@5 in Beauty, which achieves comparable scores with the best NCL baseline.
+In particular, ColaRec consistently outperforms previous CF-based and generative models in Recall@20, achieving a relative improvement of 3.87%, 3.54%, 6.72% and 3.88% on Beauty, Sports, Phone and Recipe, respectively.
 These results demonstrate the effectiveness of ColaRec and its generalization across different domains.
-Secondly, compared with DSI-R and DSI-S, which directly adapt generative retrieval methods to the recommendation task, ColaRec achieves a notable 47.89%, 38.13%, and 44.70% relative improvement in terms of Recall@5 on three datasets respectively, demonstrating effectiveness of the proposed ColaRec.
+Secondly, compared with DSI-R and DSI-S, which directly adapt generative retrieval methods to the recommendation task, ColaRec achieves a notable 47.89%, 38.13%, 80.82% and 26.11% relative improvement in terms of Recall@5 on four datasets respectively, demonstrating the effectiveness of the proposed ColaRec.
 Such results also demonstrate that naively transferring the generative retrieval methods for recommendation cannot achieve satisfying results.
-Furthermore, our method consistently outperforms the state-of-the-art generative recommendation model TIGER, which uses a RQ-VAE module to construct GIDs.
-The reason may be that TIGER only considers item content information but the collaborative signals are overlooked without an explicit alignment process.
+Furthermore, our method consistently outperforms existing representative generative methods TIGER and LC-Rec.
+For TIGER, the reason may be that it only considers item content information but the collaborative signals are overlooked without an explicit alignment process.
+While for LC-Rec, the recommendation is conducted through language modeling and generation under designed prompts, which is not fully aligned with the recommendation task, leading to sub-optimal results.
 Lastly, existing generative methods (e.g., DiffRec and TIGER), while exhibiting promising results in some specific scenarios (Rajput et al., 2023; Wang et al., 2023b; Li et al., 2023a), still underperform the strong CF-based methods (e.g., NCL) in the general recommendation task.
 In contrast, ColaRec achieves competitive results compared to these CF methods on all datasets, demonstrating the potential of infusing content information for collaborative generation of recommender systems.
+Table 3.
+Performance comparison of long-tail users.
+The best and the second-best scores are marked in bold and underlined fonts, respectively.
+** denotes the improvements are significant with p-value ¡ 0.05.
+R stands for Recall.
+CF-based Methods Generative Models for Recommendation
+Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER LC-Rec Ours
+Beauty R@5 0.0416 0.0636 0.0555 0.0639 0.0510 0.0464 0.0131 0.0415 0.0487 0.0492 0.0660**
+R@10 0.0604 0.0922 0.0825 0.0907 0.0742 0.0662 0.0228 0.0653 0.0745 0.0772 0.0975**
+R@20 0.0817 0.1253 0.1160 0.1264 0.1039 0.0917 0.0354 0.0940 0.1084 0.1107 0.1327**
+Sports R@5 0.0209 0.0433 0.0355 0.0440 0.0329 0.0267 0.0116 0.0307 0.0380 0.0397 0.0456**
+R@10 0.0317 0.0639 0.0562 0.0645 0.0495 0.0394 0.0170 0.0472 0.0581 0.0617 0.0674**
+R@20 0.0468 0.0904 0.0836 0.0908 0.0725 0.0553 0.0273 0.0728 0.0882 0.0929 0.0976**
+Phone R@5 0.0405 0.0723 0.0660 0.0727 0.0571 0.0451 0.0206 0.0404 0.0602 0.0612 0.0756**
+R@10 0.0590 0.1054 0.0986 0.1043 0.0861 0.0641 0.0371 0.0623 0.0898 0.0918 0.1131**
+R@20 0.0855 0.1482 0.1418 0.1473 0.1228 0.0899 0.0600 0.0939 0.1293 0.1353 0.1590**
+Recipe R@5 0.0128 0.0204 0.0121 0.0210 0.0182 0.0172 0.0157 0.0171 0.0181 0.0189 0.0219**
+R@10 0.0229 0.0320 0.0212 0.0322 0.0309 0.0269 0.0274 0.0295 0.0316 0.0313 0.0334**
+R@20 0.0371 0.0487 0.0343 0.0490 0.0499 0.0412 0.0443 0.0475 0.0504 0.0493 0.0528**
 5.5.2. Comparison on long-tail users
@@
 In this experiment, the ratio between head users and long-tail users is set as 20%:80%.
-Table 3 reports the results of Recall on three datasets.
+Table 3 reports the results of Recall on four datasets.
 Results of NDCG show similar trends and are omitted due to the reason of space.
@@
 The reason is that ColaRec models both user-item interactions and item content information.
-Given the long-tail users with less interaction knowledge, the content information helps ColaRec to gain better performance.
+Given the long-tail users with less interaction knowledge, ColaRec gains better performance with the help of the content information.
 To conclude, the proposed ColaRec is effective to yield better performance compared with existing baselines. This improvement is more significant on long-tail users.
+Table 4. Effect of multi-task learning and different types of GIDs. The best scores are marked in bold.
+Beauty Sports Phone Recipe
+Recall@5 NDCG@5 Recall@5 NDCG@5 Recall@5 NDCG@5 Recall@5 NDCG@5
+ColaRec 0.0667 0.0449 0.0442 0.0294 0.0745 0.0490 0.0198 0.0151
+(1) w/o textual content 0.0527 0.0346 0.0364 0.0239 0.0636 0.0426 0.0181 0.0141
+(2) w/o indexing 0.0637 0.0428 0.0422 0.0278 0.0728 0.0487 0.0179 0.0142
+(3) w/o $\mathcal{L}_{bpr}$ 0.0612 0.0412 0.0424 0.0282 0.0719 0.0486 0.0184 0.0140
+(4) w/o $\mathcal{L}_{c}$ 0.0657 0.0434 0.0422 0.0279 0.0731 0.0485 0.0188 0.0145
+(1) $iad$ 0.0658 0.0437 0.0428 0.0285 0.0719 0.0474 0.0189 0.0145
+(2) $Random$ 0.0600 0.0401 0.0411 0.0272 0.0667 0.0443 0.0190 0.0149
+(3) $Content$ 0.0662 0.0440 0.0423 0.0278 0.0716 0.0477 0.0183 0.0141
 5.6. Ablation Study (RQ2)
 In this section, we conduct ablation studies to analyse the effectiveness of each component in ColaRec.
-We implement four ablative variants of ColaRec, including: (1) w/o textual contentdeletes all textual content information in the model input; (2) w/o indexingremoves the item-item indexing task; (3) w/o $\mathcal{L}_{bpr}$removes the ranking loss $\mathcal{L}_{\text{bpr}}$; and (4) w/o $\mathcal{L}_{c}$removes the contrastive loss $\mathcal{L}_{\text{c}}$.
-The results on three datasets are shown in Table 4.
+We implement four ablative variants of ColaRec, including: (1) w/o textual contentdeletes all textual content information in the model input; (2) w/o indexingremoves the item-item indexing task; (3) w/o $\mathcal{L}_{\text{bpr}}$removes the ranking loss $\mathcal{L}_{\text{bpr}}$; and (4) w/o $\mathcal{L}_{c}$removes the contrastive loss $\mathcal{L}_{\text{c}}$.
+The results are shown in the upper part of Table 4.
 These results clearly indicate that the removal of any component from our proposed method results in a noticeable decline in overall performance.
-From the results of variant (1), we see that removing textual content information leads to a significant reduction in performance.
-Specifically, a $26.57\%$, $21.43\%$, and $17.14\%$ drop in Recall@5 is observed across the three datasets.
+From the results of variant (1), we see that removing textual content information leads to the significant performance downgrade.
+Specifically, a $20.99\%$, $17.65\%$, $14.63\%$ and $8.59\%$ drop in Recall@5 is observed across the four datasets, respectively.
 This suggests the importance of incorporating textual content information to enhance the models’ understanding of items.
-Furthermore, the variant (3), which is trained without $\mathcal{L}_{\text{bpr}}$, exhibits a notable performance decrease in comparison to ColaRec.
-This highlights the effectiveness of the pairwise ranking objective, which focuses on the relative prioritization of positive and negative items within the generative recommendation paradigm.
+Furthermore, the variant (3), which is trained without $\mathcal{L}_{\text{bpr}}$, exhibits a notable performance decrease compared to ColaRec.
+This highlights the effectiveness of the pairwise ranking objective, which focuses on the relative prioritization of positive items within the generative recommendation paradigm.
 Besides, the results of variant (2) and (4) verify the effectiveness of the proposed item-item indexing task and the contrastive objective $\mathcal{L}_{c}$.
-A consistent performance reduction is observed on the three datasets when either of the two techniques is removed.
+A consistent performance reduction is observed on the four datasets when either of the two techniques is removed.
 This suggests that aligning item content information with user-item collaborative signals not only facilitates mutual reinforcement but also enables the learning of more comprehensive and effective representations.
@@
 (b) Sports
-(c) Phone
-Figure 4. Impact of the length of GIDs. The GID length $l$ is set from $1$ to $4$.
+Figure 4. Impact of the length of GIDs.
 (a) Beauty
 (b) Sports
-(c) Phone
-Figure 5. Impact of the number of clusters. The number of clusters $K$ varies from $32$ to $128$.
+Figure 5. Impact of the number of clusters.
 5.7. GID Investigation (RQ3)
 The construction of GIDs plays a crucial role in generative recommendation as it defines the model’s search space for generation.
-To answer RQ3, we conduct two analytical experiments about GIDs: (1) an analysis of different GID types (e.g., the single token $iad$, the random assigned GID, and the content-based GID), and (2) an analysis of hyper-parameters when constructing GIDs (e.g., the length and $K$-means cluster numbers).
+To answer RQ3, we conduct two analytical experiments about GIDs: (1) an analysis of different GID types (e.g., the single token $iad$, the random assigned GID, and the content-based GID), and (2) an analysis of hyper-parameters when constructing GIDs (e.g., the length and the $K$-means cluster number).
 5.7.1. Effect of different GID types
-To further evaluate our GID generation strategy, we conducted an ablation study comparing it with three techniques: iad-based GID, $Random$ GID, and $Content$ GID.
+To further evaluate our GID generation strategy, we conduct an ablation study comparing it with three techniques: iad-based GID, $Random$ GID, and $Content$ GID.
 Specifically, the iad-based GID assigns a unique single token to each item to represent it using the corresponding vector in the item embedding matrix.
 The Random GID assigns a random string to each item as the identifier, without considering any prior knowledge.
-Content GID represents the GID constructed from item content information.
-Unlike the collaborative GID used in ColaRec, the content-based GID employs a hierarchical $K$-means clustering algorithm (Tay et al., 2022) to cluster items based on their textual representation derived from a pretrained BERT.
+The Content GID represents the GID constructed from item content information.
+Unlike the collaborative GID used in ColaRec, the content-based GID employs a hierarchical $K$-means clustering algorithm (Tay et al., 2022) to cluster items based on their textual representation derived from a pretrained BERT model.
 To make a fair comparison with ColaRec, the length and the codebook size of both $Random$ GIDs and $Content$ GIDs are identical to those in ColaRec.
-Table 5 details our results on the three datasets, where ColaRec attains the best performance.
+The bottom part of Table 4 details our results on the four datasets, where ColaRec attains the best performance.
 This indicates that our GID construction method, based on collaborative signals, is highly effective for generative recommendation.
@@
 5.7.2. Impact of GID Hyper-parameters
-In this section, we investigate the impact of hyperparameters $l$ and $K$ in the GID construction process.
+In this section, we investigate the impact of hyper-parameters $l$ and $K$ in the GID construction process.
 Figure 4 shows the results of ColaRec with different GID lengths $l$, ranging from $1$ to $4$.
-We can see that as $l$ varies, the recommedation performance exhibits some fluctuations, with the best results typically observed when $l=3$.
-It is worth noting that the GID of ColaRec consists of two parts: (i) the prefix tokens $[z_{i}^{1},\dots,z_{i}^{l-1}]$ from hierarchy clustering; and (ii) the last token $z_{l}$, which is randomly assigned to ensure uniqueness of the GID.
-A longer GID tends to include a greater proportion of hierarchy clustering parts, potentially facilitating model learning.
-However, a too long GID means more autoregressive decoding steps in model generation, which increases search difficulty and inference latency.
-Therefore, balancing efficiency and effectiveness, we have selected $l=3$ as the default setting.
-To assess how the number of clusters $K$ affects performance, we fixed the GID length $l$ as $l=3$ and varied $K$ with values $32$, $64$, $96$, and 128.
+Note that to include the whole item set, shorter GIDs indicate larger search spaces for each GID position, and vice versa.
+We can see that as $l$ varies, the recommendation performance exhibits some fluctuations.
+For the performance drop from $l=1$ to $l=2$ in Beauty and Sports, the reason is that compared with single token GIDs, GIDs with $l=2$ increase one decoding step but the search space for each step is still large, increasing the search difficulty.
+When $l=3$, ColaRec achieves a better trade-off between decoding steps and the search space for each step, leading to the best performance in most cases.
+When $l=4$, a too long GID means more autoregressive decoding steps in model generation, which increases generation difficulty and inference latency.
+Therefore, we have selected $l=3$ as the default setting in this paper.
+To assess how the number of clusters $K$ affects performance, we fix the GID length $l$ as $l=3$ and vary $K$ with values $32$, $64$, $96$, and 128.
 As shown in Figure 5, a higher $K$ typically results in a slight decrease in overall performance.
-The decrease on Beauty is more notable.
-The reason is that a higher $K$ indicates a larger search space in the decoding process, and thus increase the generation difficulty.
-To this end, choosing a suitable clustering number according to number of items is important for generative recommendation.
+The decrease in Beauty is more notable.
+The reason is that a higher $K$ indicates a larger search space in the decoding process, and thus increases the generation difficulty.
+To this end, choosing a suitable clustering number according to the number of items is important for generative recommendation.
+The principle is that $K$ needs to be large enough to encode the whole item set, after that $K$ needs to be controlled to limit the search space.
 6. Conclusions and future work
 This paper proposes ColaRec, a novel framework to conduct content-based collaborative generation for recommender systems.
-As an end-to-end generative recommender, ColaRec effectively integrates both item content information and user-item collaborative signals within a unified framework through fine-tuning an encoder-decoder based language model.
+As an end-to-end generative recommender, ColaRec effectively integrates both item content information and user-item collaborative signals within a unified framework through a generative model.
 In addition, an auxiliary item indexing task and a contrastive loss are proposed to further align the model’s representations of item content and user-item collaborative signals.
-We have conducted extensive experiments on three real-world datasets and empirical results demonstrate the effectiveness of ColaRec.
-In the future, we intend to investigate more methods to construct appropriate generative identifiers (GIDs) and adopt more effective approaches to better align content information and collaborative signals.
-We also plan to introduce larger language models together with larger volume of training data to generate better recommendation.
-Negative sampling for generative recommendation is one of our future works, either using GIDs to sample more informative negative samples or utilizing generative models to generate simulated negative instances.
-Besides, parameter efficient fine-tuning and model efficiency are also potential research directions.
+We have conducted extensive experiments on four real-world datasets and empirical results demonstrate the effectiveness of ColaRec.
+In the future, we intend to investigate more methods to construct generative identifiers (GIDs) and adopt more effective approaches to better align content information and collaborative signals.
+Negative sampling for generative recommendation is also one of our future works, either using GIDs to sample more informative negative samples or utilizing generative models to generate synthetic negative instances.
+Besides, how to improve model efficiency for generative recommendation is also a potential research direction.
+Acknowledgements
+This work was supported by the Tencent WeChat Rhino-Bird Focused Research Program (WXG-FR-2023-07), the Natural Science Foundation of China (62202271, 62372275, 62272274, T2293773,
+62102234, 62072279), the National Key R&D Program of China with grant No.2022YFC3303004, the Natural Science Foundation of Shandong Province (ZR2021QF129) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
+All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.